### v1

# Mining-Gym: A Configurable RL Benchmarking Environment for Truck Dispatch Scheduling

Abstract
Mining process optimization, particularly truck dispatch scheduling, is a critical factor in enhancing the efficiency of open-pit mining operations. However, the dynamic and stochastic nature of mining environments—characterized by uncertainties such as equipment failures, truck maintenance, and variable haul cycle times—poses significant challenges for traditional optimization methods. While Reinforcement Learning (RL) has demonstrated promise in adaptive decision-making for mining logistics, its practical deployment requires rigorous evaluation in realistic and customizable simulation environments. The lack of such standardized environment benchmarks limits fair algorithm comparisons, reproducibility, and real-world applicability of RL-based approaches in open-pit mining settings. To address this challenge, we introduce Mining-Gym, a configurable, open-source benchmarking environment designed for training, testing, and comparing RL algorithms in mining process optimization. Built on Discrete Event Simulation (DES) and seamlessly integrated with the OpenAI Gym interface, Mining-Gym offers a structured testbed that enables the direct application of advanced RL algorithms from Stable Baselines. The framework models key mining-specific uncertainties, such as equipment failures, queue congestion, and stochasticity of mining processes, ensuring a realistic and adaptive learning environment. Additionally, a graphic user interface (GUI) for easy parameter selection for mine-site configuration, comprehensive data logging system, a built-in KPI dashboard and real-time representative visualization of mine-site enables in-depth performance analysis, facilitating standardized, reproducible evaluation across multiple RL strategies and baseline heuristics.
Index Terms:
Reinforcement Learning, Truck Scheduling, Discrete Event Simulation, OpenAI Gym, Mining Simulation, Resource Allocation
I Introduction
Mining process optimization aims to enhance efficiency and productivity by improving resource allocation, equipment scheduling, and material handling. However, these operations are highly complex, influenced by dynamic factors such as equipment failures, fluctuating ore quality, and unpredictable environmental conditions. Traditional optimization methods, such as linear programming and heuristics, struggle to adapt in real time, leading to inefficiencies and increased costs.
RL offers a promising dynamic approach, but its adoption in mining remains limited due to the lack of standardized simulation environments that accurately model the stochastic nature of mining operations, provide a configurable testbed for algorithm development and comparison, and enable reproducible research with standardized scenarios and metrics. Additionally existing simulators are often proprietary, overly simplistic, or tailored to specific algorithms, hindering fair comparisons and slowing the transition from academic research to industrial application.
To address these, we introduce Mining-Gym, a configurable benchmarking framework for truck dispatch scheduling optimization. Unlike prior tools, Mining-Gym offers a comprehensive framework covering the following :
  • A high-fidelity discrete-event simulation (DES) capturing real-world complexities,
  • Seamless integration with OpenAI Gym for RL algorithm compatibility,
  • Extensive customization for diverse mining scenarios,
  • Comprehensive logging and visualization tools, and
  • Open-source availability for collaborative research.
By standardizing benchmarking, Mining-Gym enables fair algorithm evaluation, facilitates reproducible experiments, and bridges the gap between theoretical RL development and industrial application.
The rest of this paper is structured as follows: Section II reviews related work in mining simulation and RL and higlights the limitations of conventional simulators. Section III details Mining-Gym’s architecture, integrating DES with RL. Section IV covers discussions on implementation features, including the configuration system and visualizations. Section V outlines experimental setup and Section VI presents and discusses the results. Finally section VII presents the conclusion and future work.
II Background and Related Work
II-A Background
II-A1 Truck dispatch scheduling optimization
Managing dispatch scheduling, particularly truck dispatching, represents a pivotal challenge in mining process optimization. Truck dispatching is dedicated to transporting extracted supply materials—both in quantity and quality—from mining fronts, where shovels excavate, to destinations like Crushers, or dumping sites. These dispatching decisions significantly impact operational efficiency, being critically important as a large portion of mining costs are linked to truck-shovel activities Truck dispatching tasks often utilize mathematical programming to reduce equipment waiting times and optimize production [1]. Whereas heuristic methods simplify decision-making by using practical experience instead of exhaustive optimization. These strategies include assigning trucks to the nearest shovel, prioritizing by equipment capacity or material demand, and leveraging historical data patterns [2]. Truck dispatching tasks often utilize static optimization techniques to minimize equipment waiting times and optimize production [1]. However, such conventional methods typically require re-optimization when complex configurations change, such as during equipment breakdowns, and may not be robust enough to handle the stochastic nature of minesite processes.
Reference Year Uncertainties Simulator arch. RL adaptability RT Visualization Customizable Platform used Open source
[3] 2022 Extensive DES N.A. $\times$ $\checkmark$ Simpy $\times$
[4] 2022 Extensive DES N.A. Fair $\checkmark$ Flexsim $\times$
[5] 2022 Extensive DES N.A. $\times$ $\checkmark$ - $\times$
[6] 2023 Fair DES Custom $\times$ $\checkmark$ Simpy $\times$
[7] 2023 Fair Rule based Custom $\times$ $\times$ OpenAI $\times$
[8] 2023 Fair DES Custom $\times$ $\checkmark$ Python $\times$
[9] 2023 Fair Rule based N.A. $\times$ $\times$ MATLAB $\times$
[10] 2023 Extensive Rule based Custom $\times$ $\times$ - $\times$
[11] 2024 Fair DES N.A. Fair $\checkmark$ Arena $\times$
[12] 2024 Extensive DES N.A. Extensive $\checkmark$ Simpy $\checkmark$
[13] 2024 Fair DES Custom Fair $\checkmark$ DISPATCH $\times$
Ours 2025 Extensive DES Adaptable Extensive $\checkmark$ Salabim, Python $\checkmark$
TABLE I: A study of conventional simulators used in Surface Mining process optimization, especially for Truck Dispatching. Comparative analysis of mining simulators based on key features. The table evaluates simulators across several dimensions: Uncertainties, ranging from Extensive (comprehensive modeling of equipment failures, maintenance, variable haul times) to Fair (modeling of limited uncertainties); Simulator Architecture, comparing (DES) and Simpler Rule-based Logic; RL Adaptability, which indicates whether the simulator is Adaptable (native integration), Custom (requires adaptation), or N.A. (not designed for RL); Real-time Visualization, measuring the extent of Extensive (Elaborate animated visuals with KPI dashboards), Fair (basic visual representation, or None.
However, due to the limitations of static optimization methods in dynamic environments, research is increasingly focusing on alternative, adaptive approaches.
II-A2 DES and integration with RL in Mining Operations
To effectively apply RL in mining logistics, it must be integrated with robust simulation frameworks. DES has been widely used in industrial and mining applications for optimizing processes like truck scheduling. By modeling stochastic interactions between equipment and processors, DES captures variability and complexity, often replacing intricate mathematical models with probabilistic parameters [14]. Applications include supply chain evaluation [15] and ensuring adherence to production schedules [16], demonstrating DES’s role in handling uncertainty and improving operational efficiency.
Recent research has explored combining RL with DES to simulate complex industrial environments. While standard RL environments like OpenAI Gym provide useful benchmarks, they often lack the realism required for industrial applications [17]. Industrial settings demand detailed, stochastic modeling, similar to DES, to account for dynamic conditions and resource constraints. Integrating RL’s trial-and-error learning with DES has gained traction for modeling real-world stochastic systems. For example, [18] transformed DES-based SCT controllers into RL environments, enhancing decision-making in automotive plant control.
This RL-DES integration has been applied to scheduling and optimization across industries. [19] employed Deep Q-Networks (DQN) for flexible job shop problems, outperforming traditional metaheuristics, while [20] used DES and OpenAI Gym to create RL-compatible production scheduling environments, simplifying RL algorithm deployment. In mining, RL has been explored for short-term planning, truck dispatching, and scheduling. [6] introduced a curriculum-driven RL method for vehicle dispatching to address sparse rewards, while [8] developed a real-time RL-based dispatching system for autonomous trucks. Additionally, [13] applied Q-learning to optimize material supply during operational delays. These studies highlight the potential of RL-DES integration in enhancing decision-making and efficiency in mining operations.
II-A3 RL and the importance of simulator
RL excels in sequential decision-making, delivering state-of-the-art performance across various domains, including robotics, locomotion control, autonomous driving, and multi-agent systems. [21] The mining truck dispatching problem can be framed as a sequential decision-making task, where truck assignments must adapt to evolving conditions. RL provides a suitable framework for optimizing these dispatching strategies
In mining, RL enables adaptive decision-making by learning optimal dispatching policies through trial-and-error interactions. This approach accommodates dynamic changes in configurations, equipment failures, fluctuating ore quality, and weather, without frequent re-optimization. RL allows continuous refinement of dispatching policies to maximize efficiency and resilience in complex, uncertain environments. However, RL’s reliance on environmental interaction presents challenges in real mining due to safety, cost, and timeline concerns. Unlike controlled simulations real-time training is impractical as RL requires extensive exploration. Despite advancements in sample efficiency [22, 23], mining operations still require hundreds to thousands of episodes, hindering real-world deployment.
Simulators are preferred for safe, cost-effective RL training. They enable rapid learning in diverse virtual scenarios, crucial for robust policy development. Common RL simulators like OpenAI Gym [24], although simplified, form the basis for algorithm development. Simulators are vital in mining optimization due to industry complexity and variability. High-fidelity simulations improve accuracy, ensuring learned strategies transfer effectively to real-world operations, enhancing efficiency and safety, making RL a viable tool for mining logistics.
II-A4 RL Applications in Truck Dispatch Optimization
A limited number of studies have applied RL to truck dispatch optimization, but these are typically tailored to specific RL algorithms, limiting their adaptability. Adapting them to work with different RL methods would require extensive understanding and potential modifications. Huo et al. [7] apply Q-learning to optimize dispatching in haulage operations, reducing greenhouse gas emissions while maintaining production. Matsui et al. [8] develop a real-time dispatching algorithm using deep RL for autonomous haulage trucks, improving transportation efficiency and fuel consumption. De et al. [10] extend RL to short-term production planning, integrating actor-critic agents for equipment allocation and production scheduling. Chiarot et al. [13] apply Q-learning-based deep RL to reduce delays during shifts and breaks, improving material supply to crushers.
II-A5 Limitations: Existing simulators
Refer to Table I for a comparison of conventional simulators based on crucial features. A notable absence of real-time visualization is observed in several studies, including [3, 5, 6, 7, 8, 9, 10, 13], which can hinder the ability to monitor and manage mining operations effectively. Customizability is another missing feature in [7], [9], and [10], limiting the flexibility of these simulators for adapting to different mining scenarios. Furthermore, the scarcity of open-source solutions, with only [12] providing such a framework, restricts broader adoption and collaborative improvement. A number of works, such as [7], [9], and [10], utilize rule-based architectures that may not be as adaptable or robust as more advanced simulation methods. Additionally, the proprietary nature of many conventional simulators limits research replication and comparison. Many also fail to adequately account for random events affecting key mining components like trucks, shovels, or crushers, as seen in [7]. Finally, many simulators are not designed for RL settings or are tailored to specific algorithms, limiting their versatility [8].
Addressing these limitations is crucial for effectively training and testing RL algorithms to obtain optimal policies, for surface mining process optimization. Proper real-time visualization and comprehensive data logging are essential for ensuring the repeatability and comparability of experiments. Accurate simulation of real-world complexities and uncertainties, along with seamless integration with widely accepted formats like OpenAI Gym, for ease of use and training of off-the-shelf and custom algorithms.
Figure 1: Overview of the Mining-Gym framework from different perspectives. (a) Event-driven or decision-point-based RL setting: The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment. (b) Comprehensive system architecture: Displays all key components of the Mining-Gym, including the graphical user interface (GUI), the generated configuration file used to initialize the environment, and the dashboard with real-time visualizations. (c) OpenAI Gym-compatible RL interface: Illustrates how the Mining-Gym integrates with OpenAI Gym by adapting simulation signals into the standard reset and step methods. This compatibility allows seamless integration with popular RL libraries, such as Stable-Baselines3 [25], enabling easy training and testing of RL models. Note: the reward function in (a) is for demonstration only.
III Mining-Gym: System Architecture and Design
The mining process especially the dynamics and uncertainty of the Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle is captured using a DES based simulation model. While the scheduling problem is modeled as a MDP in order to solve it using RL algorithms (see Fig.1). Following we discuss these two modeling processes.
III-A Mining-Gym overview
The mining gym simulator and benchmarking tool is a sophisticated platform designed for RL applications in mining operations. It features a GUI-based interface where users input crucial parameter values that define the state of the mining site, equipment, and other relevant factors. These parameters are then translated into a human-readable and editable text file, allowing for easy adjustments directly within the simulation environment. Users can choose to either train a scheduler policy or run an existing one, which can be either classical or RL-based. The simulator includes an additional interface that supports the training and testing of modern RL algorithms, enabling advanced simulation capabilities. Furthermore, the tool provides a real-time dashboard displaying key performance indicators (KPIs) and a dynamic, visual representation of the mining site’s operations, offering users valuable insights into the system’s performance and the effects of various parameters on mining activities.
We present a DES environment for mining processes, integrated with OpenAI Gym—rebranded as Gymnasium in 2023—for RL applications. Gymnasium is a widely used toolkit that standardizes interactions across diverse environments, from control tasks to robotics and video games. It enhances the original Gym with improved modularity, better support, and expanded features, enabling efficient testing, comparison, and debugging of RL algorithms. Our wrapper maps DES inputs/outputs to Gym’s reset and step methods, ensuring compatibility with RL frameworks like Stable Baselines. Stable Baselines provides well-tested implementations of RL algorithms such as PPO, A2C, and DQN, simplifying integration with Gym environments. The done flag signals episode termination, while info offers additional simulation insights. Together, Gymnasium and Stable Baselines create a powerful ecosystem for RL research, facilitating reproducible experimentation and benchmarking in complex domains like mining simulations.
III-B Mining Process Modeling: DES Overview
For DES modeling we have used a comprehensive python package named Salabim [26], which offers process interaction methods, queue handling, resources, statistical sampling, and real-time 2D/3D animation capabilities.
Components are fundamental building blocks that define the dynamic behavior of entities within the simulation environment. By defining entities as components, SALABIM can simulate complex interactions, resource contention, and event-driven behaviors essential for realistic modeling. Trucks (lower priority) and Breakdown Events (higher priority) are modeled as components in our work. Resource is a fundamental component used to model and manage entities that are shared among components within a simulation. Resources represent facilities, equipment, or services that components (such as trucks, shovels, or processes) compete for or utilize during their activities. Specifically, in Mining-Gym, Shovels, Dumps and Crushers are modeled as resources.
Figure 2: (A) Simplified mining system showing three key components: (1) Resource Handler managing resource availability and assignments, (2) Preemption Handler detecting breakdowns and managing repair processes (B) DES-RL interaction flow illustrating how the RL policy integrates with the DES. At decision points, the environment state is processed by the RL policy to determine resource assignments. Immediate or step rewards guide learning during simulation, while the episodic reward at shift (or episode) end updates the policy before environment reset. (C) Load-Haul-Dump-Return-Query (LHDRQ) cycle illustrating the truck’s journey through the mining process, which begins with querying the dispatcher for assignments, followed by loading material, hauling to the destination, dumping, and returning empty. Breakdown events, managed by the Preemption Handler, can interrupt operations at any stage.
In the SALABIM framework, the Truck component In the SALABIM framework, the Truck component follows a Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle. In the loading phase, a truck requests an available shovel as advised by the dispatcher and waits its turn. Once granted access, it undergoes loading, simulating material transfer time. The truck then enters the haul phase, traveling a predefined trajectory to the dump (crusher or dumping site), represented by holding for travel time. The dump choice is made at the loading site or shovel.
At the dump, the truck enters the dump phase, requesting access to an available crusher or dumping site. After access, it performs dumping, mirroring the time needed to unload ore or waste. The return phase follows, where the truck travels back to the shovel, completing one haul cycle. The query phase then begins, where the truck seeks scheduling or resource allocation from the dispatcher. The Dispatcher agent is the core scheduling strategist, optimizing mining operations by balancing resource utilization and minimizing wait times. It employs basic strategies like fixed schedules or nearest-first, as well as advanced approaches such as RL-trained neural network policies.
A truck may request the following resource allocations: 1) Shovel, 2) Crusher, 3) Dumping site, and 4) Route (not currently considered). Mining-Gym’s dispatcher modules handle these requests (except routing) and supports baseline (e.g. Random) aswell as learned strategies (RL based). Only one dispatcher can be replaced with a learnable NN-based learnable policy at a time, with shovel allocation set as the default for all experiments.
Modeling breakdown events is crucial for simulating disruptions when trucks, shovels, or dumps become unavailable due to failures. These breakdowns impact loading, dumping, and transit, affecting efficiency and resource use. A preemption handler enhances realism by managing resource interruptions. For example, if a shovel breaks down mid-loading, the preemption handler pauses operations until repairs or replacement occur, accurately reflecting real-world maintenance disruptions.
Additional features : in the final Mining-Gym simulation build upon the basic workflow (Fig. 2).
  • Choice between Dump Types: The stripping ratio of waste to ore is represented by the parameter $\epsilon$, which controls the probability that a truck carrying ore is directed to the crusher rather than the dumping site. Specifically, with probability $\epsilon$, the truck is directed to the crusher, and with probability $1-\epsilon$, it is sent to the dumping site. This parameter can be adjusted to balance waste and ore management, adapting to varying operational conditions.
  • Scalable Configurations: The simulation supports adjustable configurations for Trucks, Shovels, Crushers, and Dumping sites, allowing for scalable operational simulations that can be tailored to different mining scenarios.
  • Uncertainty Modeling: To reflect the stochastic nature of real mining operations, the simulation incorporates sampling from various probability distributions. These distributions, along with their parameters, are configurable, enabling flexible modeling of operational uncertainty.
  • Dispatcher Strategies: Non-learnable, baseline dispatcher strategies (default random) for resources like Crushers and Dumping sites are included. These strategies allow for the evaluation of different dispatching approaches without relying on RL.
III-C Modeling the Dynamic Dispatching Problem as MDP
III-C1 Typical RL
In a typical RL setting, the agent interacts with the environment at every time step $t$. The environment is modeled as a Markov Decision Process (MDP) defined by the tuple $(S,A,P,R)$, where $S$ is the set of states, $A$ is the set of actions, $P(s^{\prime}|s,a)$ is the state transition probability, and $R(s,a)$ is the reward function. At each time step, the agent selects an action $a_{t}$ based on the current state $s_{t}$, receives a reward $r_{t+1}$, and transitions to a new state $s_{t+1}$. The goal is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative reward $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, where $\gamma$ is the discount factor.
III-C2 Decision-point based or Event-driven RL
In mining truck scheduling, an event-driven reinforcement learning (RL) approach is adopted, where the agent interacts with the environment only at discrete decision points $d$, rather than at every time step. A decision point represents a specific instance within the operational environment where the agent must choose an action based on the observed state $s$, which includes truck and job statuses. The reward $r$ at each decision point reflects performance metrics such as reduced waiting times and increased throughput, while the action $a$ involves dispatching trucks to different tasks. The framework aims to optimize the expected cumulative reward, given by $\mathbb{E}[\sum_{d=0}^{\infty}\gamma^{d}r_{d}]$, with agent-environment interactions constrained to these discrete decision points to align with real-world operational constraints. This approach has been widely used in complex environments where continuous interaction is impractical or unnecessary, e.g. in vehicle routing [27] and supply chain management [28].
III-C3 Defining Agent-Environment interaction and the MDP
Formulating truck dispatch scheduling in open-pit mining as a Markov Decision Process (MDP) involves defining the components of MDP: states, actions, rewards, and transitions. The Mining-Gym framework considers $TR$ trucks (represented as $\tau$) interacting with the DES based mining environment. At every decision point $d\in D$, after dumping the material into the appropriate location (crusher or dumping site), a new resource assignment is requested by a truck $\tau_{i}\;\text{where}\;i\in N$. The dispatcher agent $\mathcal{D}$ observes the current state $S_{d}\in\mathcal{S}$, where $S_{d}$ represents the current status of the mining complex’s performance at decision point $d$, and takes an action $A_{d}^{i}\in\mathcal{A}$, determining the next shovel to which the truck $i$ will be assigned.
Below, we define the various components of the MDP used in this work:
  • 1)
    States: The state represents the current status of the mining operation. The state of the system at a given time must encode all the features needed for the agent to learn a relationship with the desired objective to be maximized. For this task the state of the system is encoded as a vector with the following components
    $s^{d}=[SA_{d},TA_{d}]$ (1)
    where,
    $SA_{d}$ represents shovel-related attributes as an encoded vector at the current decision point $d$. It includes Shovel ID, encoded in 3-bit chunks, Queue Length, representing the number of trucks waiting per shovel, and Shovel Status, a binary indicator of shovel availability (online/offline). Similarly, $TA_{d}$ captures truck-related attributes, including TruckID, encoded in 5-bit chunks, Trips complete, tracking the number of trips per truck, and Trip Status, a multi-bit representation of the truck’s current state, such as loading, transit, or maintenance (see Table: II).
    Category Attribute / Status Description
    Shovels $(SA_{d})$ Shovel ID 3-bit binary
    Queue Length Normalized float
    Shovel Status Binary (1 or 0)
    Trucks $(TA_{d})$ Truck ID 5-bit binary
    Trips Complete Normalized float
    Truck Status 3-bit binary
    Truck Status Codes
    Truck Status Binary Code
    At Shovel 000
    At Crusher 001
    At Dumping Site 010
    Moving from Shovel to Crusher 011
    Moving from Shovel to Dumping Site 100
    Moving from Crusher to Shovel 101
    Moving from Dumping Site to Shovel 110
    Breakdown 111
    TABLE II: State Space Description and Truck Status Codes
  • 2)
    Actions: Actions are decisions made at each decision step (d), by the RL policy.
    In our framework, an action corresponds to resource or shovel allocation and the action space is defined as:
    $A_{d}\in\{1,2,\dots,SH\}$
    where $SH$ represents the total number of shovels, and the action value indicates the shovel ID to which a truck is assigned at decision point $d$.
  • 3)
    Rewards: The reward function quantifies both immediate and cumulative benefits, defining objectives to maximize. In our event-based RL setting, shifts are episodes with sparse rewards given for critical milestones like meeting production targets. The environment features a long-horizon episodic reward and intermediate global rewards to guide learning.
    Once the dispatching agent $\mathcal{D}$ outputs action $A_{i}^{d}$, the DES environment responds with:
    $R_{i}^{d}=r_{\text{imm}}^{d}+\big{(}r_{\text{epi}}^{d}\;\text{if}\;d=d_{T}\;\text{else}\;0\big{)}$ (2)
    where $r_{\text{epi}}^{d}$ is the episodic reward at the end of an episode (or shift), and $r_{\text{imm}}^{d}$ is the immediate reward per decision-step $d$.
    Immediate Reward Formulation: The immediate reward uses an exponentially weighted sliding window over the latest $k$ decision points:
    $r_{\text{imm}}^{d}$ $=-\alpha\cdot\hat{TT}_{\text{Avg}}-\beta\cdot\hat{Q}_{\text{Avg}_{d}}-\gamma\cdot(1-D_{\text{DivScr}})$ (3)
    $=-\alpha\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot\tau_{j}}{\sum_{j=i-k+1}^{i}w_{j}}-\beta\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot Q_{{SH}_{j}}}{\sum_{j=i-k+1}^{i}w_{j}}$
    $\quad-\gamma\cdot(1-D_{\text{DivScr}})$ (4)
    where $w_{j}=e^{\lambda(j-i+k)}$ represents exponential weighting, and $\alpha,\beta,\gamma$ are weighting parameters such that $\alpha+\beta+\gamma=1$.
    The components are:
    Average Trip Time:
    $\tau_{d}=\frac{1}{|\Omega(d)|}\sum_{\omega\in\Omega(d)}\mathcal{T}_{\omega}(d)$ (5)
    Shovel Queue Time:
    $Q_{{SH}_{d}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\frac{1}{|\Omega_{\gamma}(d)|}\sum_{\omega\in\Omega_{\gamma}(d)}Q_{\omega}(d)$ (6)
    Diversity Score:
    $D_{\text{DivScr}}=\frac{|\{\gamma\in\Gamma\mid A_{\gamma}>0\}|}{|\Gamma|}$ (7)
    where $A_{\gamma}$ is the number of trucks assigned to shovel $\gamma$ in the recent window.
    Episodic Reward Formulation: The episodic reward balances production efficiency and resource utilization:
    $r_{\text{epi}}^{d_{T}}$ $=\omega_{1}\cdot P_{\text{Efficiency}}-\omega_{2}\cdot D_{\text{Imbalance}}$ (8)
    $=\omega_{1}\cdot\min\left(1,\frac{P_{\text{Vol}}}{P_{\text{Vol}_{\text{Target}}}}\right)-\omega_{2}\cdot\frac{D_{\text{Penal}}}{\sum_{\gamma\in\Gamma}A_{\gamma}}$ (9)
    where: $P_{\text{Vol}}$ is total material produced per shift, $P_{\text{Vol}_{\text{Target}}}$ is the production target per shift. $D_{\text{Penal}}=\max_{\gamma\in\Gamma}A_{\gamma}-\min_{\gamma\in\Gamma}A_{\gamma}$ measures shovel utilization imbalance, $\omega_{1},\omega_{2}$ are weighting parameters where $\omega_{1}+\omega_{2}=1$.
    This structure balances short-term efficiency with long-term production goals, guiding RL agents in open-pit mining operations. During each episode (e.g., shifts of 12 hours), the agent takes $N_{\text{steps}}$ actions, and the discounted sum of rewards defines the return $G_{t}$:
    $G_{t}=\sum_{k=t+1}^{N_{\text{steps}}}\gamma^{k-t-1}R_{k}$
    where $\gamma$ is the discounting factor that determines the impact of future actions on the objective function. The objective is to maximize the expected return $G_{t}$ by training the agent to improve actions, ensuring trucks meet production targets and minimize queue formation:
    $\max_{\pi}\mathbb{E}[G_{t}|S_{t}=s_{t},A_{t}=a_{t}]$
    Policy: The policy defines the strategy that the agent follows to select actions (shovel IDs) based on the current state of the system. Given the state $s_{d}=[SA_{d},TA_{d}]$, which encodes the current shovel and truck attributes, the policy predicts the shovel ID to which a truck should be assigned at decision point $d$. The policy maximizes cumulative reward by optimizing truck-shovel assignments for operational efficiency, adapting over time based on rewards to meet short- and long-term production goals, ensuring efficient dispatch and shovel utilization. The policy $\pi_{\theta}(a|s)$ is a function that gives the probability of taking action $a$ (i.e., assigning a specific shovel ID) given the current state $s$:
    $\pi_{\theta}(a|s)=\mathbb{P}(A_{d}=a|S_{d}=s)$
    where $A_{d}$ is the action taken at decision point $d$, which is the shovel ID assigned to a truck, and $S_{d}=s$ is the state of the system at decision point $d$, including both shovel-related attributes $S_{A_{d}}$ and truck-related attributes $T_{A_{d}}$. This ensures that the agent’s actions evolve to maximize long-term efficiency and achieve the desired operational goals.
IV Implementation Features
In this section we introduce different visualizations and other GUI-based tools that we provide with the MINING-GYM simulator, for better performance comparison and studying change in KPIs
IV-A Configuration System
Variables Source
A. Operational Parameters
Number of Equipment: -
Num. of Trucks TR User
Num. of Shovels SH User
Num. of Crushers CR User
Num. of Dumps DS User
Queue Size (waiting for resource) $Q$: -
At Shovel, Crusher, Dumping site $Q_{SH},Q_{CR},Q_{DS}$ DES
Load per trip $L$
Truck carrying Waste, Ore $L_{W},L_{O}$ User
Truck Speed: -
Empty truck, Loaded truck $S_{\text{EmTR}},S_{\text{LodTR}}$ User
Others: -
Number of trips (SH-CR-DS-SH) $N$ DES
Available (operation unit) time $T_{\text{SHF}}$ User
Dumping and maneuver time (min) $T_{\text{DM}}$ User
Shift duration in minutes $S_{\text{dur}}$ User
Num_shifts SN User
B. Cost and Financial Parameters
Known cost, Estimated cost $C_{\text{KW}},C_{\text{EST}}$ User
C. Equipment Performance and Efficiency
Unit Fuel Consumption ($F_{\text{Unit}}$): -
Truck with Waste, Ore, Empty $F_{W},F_{O},F_{E}$ User
Equipment off time $OF_{E}$
Truck, Shovel, Crusher, Dumping Site $OF_{\text{TR}},OF_{\text{SH}},$ DES
$OF_{\text{CR}},OF_{\text{DS}}$
Equipment idle time $IDL_{E}$
Truck, Shovel, Crusher, Dumping Site $IDL_{\text{TR}},IDL_{\text{SH}},$ DES
$IDL_{\text{CR}},IDL_{\text{DS}}$
Equipment mean time between failure: -
Shovel, Truck, Crusher, Dumping Site $F_{\text{SH}},F_{\text{TR}},F_{\text{CR}},F_{\text{DS}}$ User
Equipment mean time to repair: -
Shovel, Truck, Crusher, Dumping Site $R_{\text{SH}},R_{\text{TR}},R_{\text{CR}},R_{\text{DS}}$ User
D. Loading and Other Time:
Truck loading TRL User
Dump to Shovel DTS User
Shovel to Dump STD User
Truck dumping @ Dump TRDM User
Crusher to Shovel CTS User
Shovel to Crusher STC User
Truck dumping @ Crusher TRCR User
E. Other Parameters:
Probability of choosing Crusher $\epsilon$ User
Target Production volume $P_{\text{Vol, Targ}}$ User
TABLE III: Overall list of variables and parameters used
Figure 3: (a) Shows a sample page from the GUI interface and the resulting configuration file (b) Shows the Real-time representative visualization of the mine-site. The screenshot displays trucks queued at shovels, with the rightmost shovel offline. It also shows trucks moving between the dumping site, crushers, and the shovels.
As discussed in the previous section, a large number of parameters must be initialized to run the simulator (see Table III and Table V). To provide an organized method for supplying this information, we have designed a GUI interface. This interface allows users to set values and distributions for stochastic parameters, which are then saved to a human-readable text file. This configuration file is used by the simulator to initialize its parameters during operation (see Fig. 3(a)).
IV-B Other Visualization Features
We provide a real-time representative visualization of the mining setup as configured in the setup file. It features animated truck movements between loading points, crushers, and dumping sites, along with real-time queue formation. The visualization is runnable both during training and when executing a trained model or any other algorithm (see Fig. 3(b)).
Key Performance Indicators (KPIs) in mining process optimization include metrics like equipment utilization, cycle time, production throughput, fuel efficiency, and downtime. These metrics help evaluate the efficiency, productivity, and cost-effectiveness of mining operations, guiding improvements and strategic decisions to enhance overall performance and profitability. A key factor in KPI measurement is the time frame, which can range from shifts to daily, weekly, monthly, or yearly intervals. By default, we consider a 12-hour shift as the standard, but this can be customized as a parameter.
We include the following essential KPIs :
  1. 1.
    Total Production ($P_{VOL}$): Refers to the total quantity of material (ore + waste) excavated per shift. It is measured in tons, indicating the scale and efficiency of production activities.
    $P_{VOL}$ $=N\times L$
    where: $N$ = Total number of trips in a shift, L = Material load per trip (tons)
  2. 2.
    Trips per Hour ($\text{TPH}_{t}$): The “Trips per Hour” metric quantifies the rate at which the full haul cycle (SH-CR-DS-SH) is completed, normalized to an hourly basis. It provides a standardized measure of fleet productivity by capturing the number of new trips executed within the most recent hourly interval.
    $\text{TPH}(t)=\frac{\Delta t}{60}\left[N(t)-N(t-\Delta t)\right]$
    where $N(t)$ = Total number of trips (SH-CR-DS-SH) completed by all trucks up to time $t$, $N(t-\Delta t)$ = Total trips at $t-\Delta t$, $\Delta t$ = Measurement interval in minutes (typically 60).
    Total Number of Trips:
    $N(t)=\sum_{i=1}^{TR}C_{i}(t)$
    where $TR$ = Total number of trucks in the fleet, $C_{i}(t)$ = Trips completed by truck $i$ up to time $t$.
    Figure 4: Conceptual Diagram of Minesite
  3. 3.
    Cost per Ton ($CPT$): Refers to the financial expenditure incurred for every unit weight (typically one ton) of material produced, processed, or transported.
    $CPT$ $=\frac{C_{KW}+C_{EST}}{P_{VOL}}$
    where $C_{KW}$ = Total known cost and $C_{EST}$ = Total estimated cost
  4. 4.
    Fuel Consumption ($FC$): measures the amount of fuel consumed by mining equipment or vehicles relative to the total material hauled during operations:
    $FC=\frac{\text{Total fuel consumed}}{P_{\text{Vol}}}$
    where the Total fuel $\text{Total fuel consumed}=N\times F_{\text{Unit}}$, and the $F_{\text{Unit}}=F_{\text{Unit}}=F_{W}+F_{O}+F_{E}.$
    Here, $N$ is the number of trips completed by all trucks, $F_{W}$, $F_{O}$, and $F_{E}$ are the fuel consumptions for waste, ore, and empty trips, respectively, and $P_{\text{Vol}}$ is the target production volume.
We have also provided an interactive dashboard as part of this benchmark to analyze change in these KPIs in real time.
V Experimental Setup
V-1 Environment Design
We consider a medium sized Bauxite ($5-10$ Million Tons output) mine with 3 Shovels, 20 Trucks, 4 Crushers and 2 Dumping sites, detailed specifications are presented in Table: IV and V. A conceptual diagram of the minesite is provide in Fig. 4.
Equipment Model/ Make Quantity Specification (each)
Shovels per pit CAT 6040 2 (total 4) 43.7 t bucket size
Trucks CAT 777 20 98.2 nominal payload
Crushers Metso Lokotrack (LT106) 3 450 t / hr capacity
Dumping site - 2 Unlimited
TABLE IV: Equipment List with Specifications
V-2 Training Setup
To ensure that the trained agent generalizes well across the testing scenarios, we not only incorporate stochastic parameters to randomize breakdown and repair times for mining equipment and trucks but also enhance the reward function by introducing queue buildup metrics (queue length) and diversity score to prevent repeatedly selecting same shovels.
The parameter values used during training is provided in Table V. For training an RL policy we have used the PPO [29]algorithm from StableBaselines3 [25]). We have used all the default parameter and training settings as present in the repository’s PPO implementation. Note the scenario parameters which help specify the failure scenarios to be simulated during training or testing. For this proof-of-concept implementation and demonstration the results of only $400$ episodes or shifts of training.
Figure 5: Comparison of “Trips per hour” over the shift under different failure scenarios: (a) No failure, (b–d) $(SH_{0},TR_{0})$ values where $SH_{0}$ (Num. Shovels failed) and $TR_{0}$ (Num. Trucks failed) are set as follows—(b) $(0,6)$, (c) $(1,0)$, (d) $(1,6)$. Solid lines represent mean values, and shaded regions indicate variance from repeated runs. For shovels, failures initiate between 100–300 minutes ( 2–5 hours), and for trucks, between 100–500 minutes ( 2–8 hours) within a 12-hour shift.
V-3 Testing Setup
We designed scenarios to test the efficacy of the RL trained scheduler against a vanilla scheduler (random scheduler) as baseline, under different challenging scenarios.
For a random scheduler, the dispatcher’s action $A_{d}$ is selected randomly from the available set of shovel IDs $\{1,2,\dots,SH\}$. The selection follows a uniform distribution,
$P(A_{d}=a|S_{d}=s)=\frac{1}{SH}\quad\text{for all }a\in\{1,2,\dots,SH\}$
i.e. any state $S_{d}=s$, each action $A_{d}=a$ (where $a$ is a shovel ID) is equally likely to be chosen, with a probability of $\frac{1}{SH}$. In contrast, the RL policy $\pi_{\theta}(a|s)$ is a learned policy, where the action $A_{d}=a$ is selected based on the current state $S_{d}=s$ and the parameters $\theta$. The policy is defined as:
$\pi_{\theta}(a|s)=P(A_{d}=a|S_{d}=s)$
This probability is learned through RL algorithm where the agent adjusts $\theta$ based on rewards from its actions, optimizing its decision-making process over time. We tested the schedulers under different challenging scenarios:
  1. 1.
    All equipment available, represented as $(SH_{0},TR_{0})=(0,0)$, i.e., no Shovels and Trucks offline.
  2. 2.
    Six Trucks under maintenance, i.e., $(SH_{0},TR_{0})=(0,6)$.
  3. 3.
    One Shovel offline, i.e., $(SH_{0},TR_{0})=(1,0)$.
  4. 4.
    One Shovel and Six Trucks unavailable, i.e., $(SH_{0},TR_{0})=(1,6)$.
All failure scenarios are repeated according to a consistent temporal schedule over five repeat trials, with the inherent stochasticity in event timings due to probabilistic distributions. These repeats are then considered in the results section for analysis and discussion.
Stochastic Parameters Probability Distributions
Fuel consumption Liters
Loaded Truck (FL) in lt/ton-km $\text{Normal}(0.25,0.001)$
Empty Truck (FE) $\text{Normal}(0.35,0.001)$
Shovel (Lt/hr) $\text{Uniform}(250,300)$
Equipment specifications Value
Truck payload (TP) $\text{Normal}(98,0.5)$ t
Shovel bucket capacity (SCP) $\text{Normal}(43,0.5)$ t
Empty Truck Speed (TS) $\text{Normal}(40,60)$ km/hr
Loaded Truck Speed (TL) $\text{Normal}(25,40)$ km/hr
Equipment Maintenance** Time (in mins )
MTBF Shovel (FSH) $\text{Poisson}(2520)$
MTBF Truck (FTR) $\text{Poisson}(2160)$
MTBF Crusher (FCR) $\text{Poisson}(30000)$
MTBF Dumping Site (FDS) $\text{Poisson}(60000)$
MTTR Shovel (RSH) $\text{Poisson}(240)$
MTTR Truck (RTR) $\text{Poisson}(300)$
MTTR Crusher (RCR) $\text{Poisson}(240)$
MTTR Dumping Site (RDS) $\text{Poisson}(120)$
Loading and Maneuver Time Time (in mins.)
Time Truck Loading (TRL) $\text{Normal}(8,2)$
Travel Time Shovel to Dump (STD) $\text{Normal}(15,5)$
Time Truck Dumping (TRDM) $\text{Normal}(5,1)$
Travel Time Dump to Shovel (DTS) $\text{Normal}(15,5)$
Travel Time Shovel to Crusher (STC) $\text{Normal}(15,2)$
Time Truck Unloading at Crusher (TRCR) $\text{Normal}(5,1)$
Travel Time Crusher to Shovel (CTS) $\text{Normal}(15,2)$
Scenario Parameters -
Shovel To Fail (STF) Binomial (2, 0.5)
Trucks To Fail (TTF) Binomial (6, 0.5)
Shovel Initial Breakdown (SIB) Uniform (100,300)
Truck Initial Breakdown (TIB) Uniform (100,500)
TABLE V: Distributions for stochastic parameters used in the example. MTBF: Mean Time Between Failures, MTTR: Mean Time To Repair.
Figure 6: Comparison of production volume $P_{vol}$ in metric tonnes per shift, averaged over 5 repeats trials.
VI Results and Discussion
The RL scheduler demonstrated significant performance improvements over the random scheduler across all scenarios, see Fig.5 and Fig. 6 :
  • Trips Per Hour: RL increased productivity by 18.2% on average, with the greatest improvement (27.2%) observed in the most challenging scenario (1 shovel and 6 trucks offline).
  • Production Volume: RL delivered 19.7% higher production volumes on average, with improvements ranging from 14.3% to 23.3% across all scenarios.
The performance advantage of RL was most pronounced in scenarios with multiple resource constraints, suggesting that RL’s optimization capabilities become more valuable as operational complexity increases. As is showing through the gain in production volume (see Fig.6) RL improved production volume across different scenarios. In the (0,0) case, where all equipment was available, the increase was 18.4%. With six trucks offline (0,6), the improvement rose to 22.7%. When one shovel was offline (1,0), production volume increased by 14.3%. Under both constraints (1,6), RL achieved the highest improvement at 23.3%.
The RL scheduling (RL-PPO) consistently outperforms random scheduling baseline method across multiple metrics and scenarios, maintaining higher productivity, especially mid-shift. Its intelligent decisions optimize resource use, making it highly effective in real-world mining operations with equipment constraints, leading to increased production.
VII Conclusion and Future Work
The introduction of Mining-Gym addresses a critical gap in mining process optimization by providing a configurable, open-source benchmarking environment for RL algorithm evaluation in truck dispatch scheduling. Our framework offers high-fidelity DES modeling of real-world complexities, seamless integration with OpenAI Gym, and comprehensive visualization tools. Preliminary experiments show that RL-based scheduling achieves up to 27.2% higher productivity and 23.3% greater production volume, with the most significant gains in resource-constrained scenarios. Mining-Gym provides a standardized evaluation framework, ensuring reproducible research and supporting the adoption of RL solutions in industrial mining operations.
The development of Mining-Gym as a benchmarking environment will continue in the following directions:
  • Enhanced Simulation Fidelity: Future iterations will improve traffic modeling, incorporate geographical factors, model heterogeneous equipment fleets, and include additional operational constraints to enhance simulation realism.
  • Expanded Scenario Library: We will develop a broader set of standardized test scenarios to cover diverse challenges such as weather disruptions, varying ore characteristics, and dynamic production targets.
  • Multi-objective Optimization Benchmarks: The framework will be extended to support multi-objective optimization, balancing production, cost, equipment lifespan, and environmental impacts—reflecting real-world mining decision-making.
  • Integration with Digital Twin Capabilities: Enhancing Mining-Gym with digital twin functionalities will connect simulations with real-world data, facilitating data-driven optimizations and bridging the gap between simulation and operational implementation.
Acknowledgments
This research was partly supported by the Advance Queensland Industry Research Fellowship AQIRF024-2021RD4.

### v2

@@
-Mining-Gym: A Configurable RL Benchmarking Environment for Truck Dispatch Scheduling
+Mining-Gym: A Configurable RL Benchmarking Environment for Open-Pit Truck Dispatch Optimization
 Abstract
-Mining process optimization, particularly truck dispatch scheduling, is a critical factor in enhancing the efficiency of open-pit mining operations.
-However, the dynamic and stochastic nature of mining environments—characterized by uncertainties such as equipment failures, truck maintenance, and variable haul cycle times—poses significant challenges for traditional optimization methods.
-While Reinforcement Learning (RL) has demonstrated promise in adaptive decision-making for mining logistics, its practical deployment requires rigorous evaluation in realistic and customizable simulation environments.
-The lack of such standardized environment benchmarks limits fair algorithm comparisons, reproducibility, and real-world applicability of RL-based approaches in open-pit mining settings.
-To address this challenge, we introduce Mining-Gym, a configurable, open-source benchmarking environment designed for training, testing, and comparing RL algorithms in mining process optimization.
-Built on Discrete Event Simulation (DES) and seamlessly integrated with the OpenAI Gym interface, Mining-Gym offers a structured testbed that enables the direct application of advanced RL algorithms from Stable Baselines.
-The framework models key mining-specific uncertainties, such as equipment failures, queue congestion, and stochasticity of mining processes, ensuring a realistic and adaptive learning environment.
-Additionally, a graphic user interface (GUI) for easy parameter selection for mine-site configuration, comprehensive data logging system, a built-in KPI dashboard and real-time representative visualization of mine-site enables in-depth performance analysis, facilitating standardized, reproducible evaluation across multiple RL strategies and baseline heuristics.
-Index Terms:
+Optimizing the mining process—particularly truck dispatch scheduling—is a key driver of efficiency in open-pit operations.
+However, the dynamic and stochastic nature of these environments, marked by uncertainties such as equipment failures, truck maintenance, and variable haul cycle times, poses significant challenges for traditional optimization.
+While Reinforcement Learning (RL) shows strong potential for adaptive decision-making in mining logistics, practical deployment hinges on thorough evaluation in realistic, customizable simulation environments.
+Yet the absence of standardized benchmarking currently hampers fair algorithm comparison, reproducibility of results, and real-world applicability of RL-based solutions in open-pit mining.
+To bridge this gap, we present Mining-Gym—a configurable, open-source benchmarking environment for training, testing, and evaluating RL algorithms in mining process optimization.
+Mining-Gym is built on Salabim-based Discrete Event Simulation (DES) and integrates with Gymnasium (formerly OpenAI Gym), enabling seamless adoption of state-of-the-art RL algorithms from Stable Baselines3.
+The framework captures mining-specific uncertainties such as equipment failures, queue congestion, and stochastic behavior through an event-driven decision-point architecture.
+Mining-Gym provides a graphical user interface (GUI) for intuitive parameter configuration, extensive data logging, and real-time visualization—supporting thorough performance analysis and standardized, reproducible evaluation of diverse RL strategies and heuristic baselines.
+To validate benchmarking performance, we compare classical heuristics with RL-based scheduling across six scenarios ranging from normal operation to severe equipment failures.
+The results demonstrate that Mining-Gym is an effective, reproducible testbed for mining-logistics research—enabling fair evaluation of adaptive decision-making for real mines.
+The demonstrative experiments also show the strong performance potential of RL-trained schedulers, motivating further development in this direction.
+keywords:
 Reinforcement Learning, Truck Scheduling, Discrete Event Simulation, OpenAI Gym, Mining Simulation, Resource Allocation
-I Introduction
+[…]
+[inst1]organization=School of Electrical Engineering and Robotics, Queensland University of Technology,addressline=2 George Street, city=Brisbane, postcode=4000, state=QLD, country=Australia
+1 Introduction
 Mining process optimization aims to enhance efficiency and productivity by improving resource allocation, equipment scheduling, and material handling.
-However, these operations are highly complex, influenced by dynamic factors such as equipment failures, fluctuating ore quality, and unpredictable environmental conditions.
-Traditional optimization methods, such as linear programming and heuristics, struggle to adapt in real time, leading to inefficiencies and increased costs.
-RL offers a promising dynamic approach, but its adoption in mining remains limited due to the lack of standardized simulation environments that accurately model the stochastic nature of mining operations, provide a configurable testbed for algorithm development and comparison, and enable reproducible research with standardized scenarios and metrics.
-Additionally existing simulators are often proprietary, overly simplistic, or tailored to specific algorithms, hindering fair comparisons and slowing the transition from academic research to industrial application.
-To address these, we introduce Mining-Gym, a configurable benchmarking framework for truck dispatch scheduling optimization.
-Unlike prior tools, Mining-Gym offers a comprehensive framework covering the following :
-  • A high-fidelity discrete-event simulation (DES) capturing real-world complexities,
-  • Seamless integration with OpenAI Gym for RL algorithm compatibility,
-  • Extensive customization for diverse mining scenarios,
-  • Comprehensive logging and visualization tools, and
-  • Open-source availability for collaborative research.
-By standardizing benchmarking, Mining-Gym enables fair algorithm evaluation, facilitates reproducible experiments, and bridges the gap between theoretical RL development and industrial application.
-The rest of this paper is structured as follows: Section II reviews related work in mining simulation and RL and higlights the limitations of conventional simulators.
-Section III details Mining-Gym’s architecture, integrating DES with RL.
-Section IV covers discussions on implementation features, including the configuration system and visualizations.
-Section V outlines experimental setup and Section VI presents and discusses the results.
-Finally section VII presents the conclusion and future work.
-II Background and Related Work
-II-A Background
-II-A1 Truck dispatch scheduling optimization
-Managing dispatch scheduling, particularly truck dispatching, represents a pivotal challenge in mining process optimization.
-Truck dispatching is dedicated to transporting extracted supply materials—both in quantity and quality—from mining fronts, where shovels excavate, to destinations like Crushers, or dumping sites.
-These dispatching decisions significantly impact operational efficiency, being critically important as a large portion of mining costs are linked to truck-shovel activities Truck dispatching tasks often utilize mathematical programming to reduce equipment waiting times and optimize production [1].
-Whereas heuristic methods simplify decision-making by using practical experience instead of exhaustive optimization.
-These strategies include assigning trucks to the nearest shovel, prioritizing by equipment capacity or material demand, and leveraging historical data patterns [2].
-Truck dispatching tasks often utilize static optimization techniques to minimize equipment waiting times and optimize production [1].
-However, such conventional methods typically require re-optimization when complex configurations change, such as during equipment breakdowns, and may not be robust enough to handle the stochastic nature of minesite processes.
-Reference Year Uncertainties Simulator arch. RL adaptability RT Visualization Customizable Platform used Open source
+However, these operations are inherently complex and uncertain, influenced by dynamic and stochastic factors such as equipment failures, fluctuating ore quality, and unpredictable environmental conditions.
+These characteristics introduce significant levels of imprecision and variability into decision-making, making it challenging for traditional optimization methods—such as linear programming and heuristics—to adapt effectively in real time.
+As a result, these conventional approaches often lead to inefficiencies and increased operational costs [1].
+Reinforcement Learning (RL), a branch of machine learning grounded in soft computing principles, offers a promising adaptive approach for handling such complexity.
+However, its adoption in mining remains limited, primarily due to the absence of standardized benchmarking environments that enable fair algorithmic comparison, reproducible experimentation, and systematic evaluation across diverse operational scenarios.
+Moreover, existing simulators are often proprietary [2], overly simplistic [3], or narrowly tailored to specific algorithms—factors that hinder fair comparisons and slow the transition from academic development to industrial application.
+To bridge this gap, we present Mining-Gym, a configurable benchmarking framework for truck dispatch scheduling optimization, designed specifically to support standardized evaluation of RL and classical algorithms in mining contexts.
+Unlike prior tools, Mining-Gym provides a comprehensive and open-source platform that supports the development and evaluation of RL algorithms by incorporating:
+  • 1.
+    A high-fidelity discrete-event simulation (DES) capturing real-world operational complexities and stochastic behaviors,
+  • 2.
+    Seamless integration with OpenAI Gym to ensure compatibility with modern RL algorithms,
+  • 3.
+    Extensive customization options to model diverse and realistic mining scenarios,
+  • 4.
+    A comprehensive system for logging, visualization, and performance tracking, and
+  • 5.
+    Open-source  [^1] Source code available at https://github.com/CBCodeHub/Mining-Gym availability to promote transparency, collaboration, and reproducibility in research.
+By enabling standardized benchmarking and fostering robust comparisons across algorithms and heuristics, Mining-Gym supports the practical application of RL in mining logistics.
+It serves as a crucial bridge between theoretical research in soft computing and its deployment in real-world industrial systems.
+The framework’s primary contribution is providing researchers with a reliable testbed for rigorous algorithmic evaluation.
+As a proof-of-concept demonstration, we present comparative results between classical heuristics (random, fixed) and RL-based scheduling (PPO) across six operational scenarios, validating Mining-Gym’s effectiveness as a benchmarking platform.
+The remainder of this paper is organized as follows: Section 2 reviews related work in mining simulation and reinforcement learning, emphasizing the limitations of current benchmarking approaches.
+Section 3 describes the architecture of Mining-Gym, which integrates discrete-event simulation with RL to facilitate standardized evaluation.
+Section 4 presents the implementation details, including the configuration system, visualization tools, and logging mechanisms that ensure reproducible experiments.
+Section 5 outlines the experimental setup and provides a demonstrative comparison between classical heuristics and RL-based scheduling to showcase the framework’s benchmarking capabilities.
+Section 6 discusses the results of these experiments.
+Finally, Section 7 concludes the paper and highlights directions for future work.
+2 Background and Related Work
+2.1 Background
+2.1.1 Truck dispatch scheduling optimization
+Open-pit mining extracts ore through a coordinated process: hydraulic shovels excavate ore and waste at the pit face, trucks transport material to processing facilities, crushers process the ore, and dumping sites stockpile waste.
+The dispatch problem focuses on optimizing truck-to-shovel assignments to minimize idle time while respecting equipment constraints.
+Key operational features in the mining setting are:
+  • 1.
+    Multi-destination routing: Ore is routed to crushers and waste to dumping sites; the dispatcher must balance this allocation.
+  • 2.
+    Equipment capacity constraints: Shovels and crushers have finite capacities, leading to potential queues and bottlenecks.
+  • 3.
+    Stochastic disruptions: Equipment failures and variable cycle times require adaptive, real-time decision-making.
+  • 4.
+    Cost drivers: Truck idle time dominates operational costs; minimizing queue congestion is therefore economically critical.
+Dispatching decisions significantly influence operational efficiency, as a large portion of mining costs are linked to truck–shovel activities.
+Truck dispatching tasks often employ mathematical programming approaches to reduce equipment waiting times and optimize production [4].
+Heuristic methods are widely used for their simplicity and flexibility, especially in dynamic mining environments where traditional optimization struggles with real-time disruptions like equipment failures or changing conditions [5].
+Common strategies include assigning trucks to the nearest shovel, prioritizing by equipment capacity or material demand, and using historical data trends.
+However, they lack guarantees of global optimality and may underperform in complex or unfamiliar scenarios.
+In contrast, static optimization approaches are often applied to truck dispatching tasks to minimize equipment idle time and enhance production efficiency [4].
+However, these traditional methods usually require full re-optimization when faced with complex changes—like equipment breakdowns—and are often not robust enough to manage the inherent uncertainty of mining operations.
+Given the limitations of static optimization methods in handling dynamic and uncertain mining environments, there is a growing research emphasis on developing adaptive and flexible approaches that can respond effectively to real-time operational disruptions.
+2.1.2 DES-RL integration in Mining Operations
+To effectively apply RL in mining logistics, it must be integrated with robust simulation frameworks.
+DES has been widely used in industrial and mining applications for optimizing processes like truck scheduling.
+By modeling stochastic interactions between equipment and processors, DES captures variability and complexity, often replacing intricate mathematical models with probabilistic parameters [6].
+Applications include supply chain evaluation [7] and ensuring adherence to production schedules [8], demonstrating DES’s role in handling uncertainty and improving operational efficiency.
+Recent research has explored combining RL with DES to simulate complex industrial environments.
+While standard RL environments like OpenAI Gym provide useful benchmarks, they often lack the realism required for industrial applications [9].
+Industrial settings demand detailed, stochastic modeling, similar to DES, to account for dynamic conditions and resource constraints.
+Integrating RL’s trial-and-error learning with DES has gained traction for modeling real-world stochastic systems.
+For example, [10] transformed DES-based SCT (supervisory control theory) controllers into RL environments, enhancing decision-making in automotive plant control.
+This RL-DES integration has been applied to scheduling and optimization across industries.
+[11] employed Deep Q-Networks (DQN) for flexible job shop problems, outperforming traditional metaheuristics, while [12] used DES and OpenAI Gym to create RL-compatible production scheduling environments, simplifying RL algorithm deployment.
+In mining, RL has been explored for short-term planning, truck dispatching, and scheduling.
+[13] introduced a curriculum-driven RL method for vehicle dispatching to address sparse rewards, while [14] developed a real-time RL-based dispatching system for autonomous trucks.
+Additionally, [15] applied Q-learning to optimize material supply during operational delays.
+These studies highlight the potential of RL-DES integration in enhancing decision-making and efficiency in mining operations.
+2.1.3 RL and the importance of simulator
+RL excels in sequential decision-making, delivering state-of-the-art performance across various domains, including robotics, locomotion control, autonomous driving, and multi-agent systems [16].
+The mining truck dispatching problem can be framed as a sequential decision-making task, where truck assignments must adapt to evolving conditions.
+RL provides a suitable framework for optimizing these dispatching strategies
+In mining, RL enables adaptive decision-making by learning optimal dispatching policies through trial-and-error interactions.
+This approach accommodates dynamic changes in configurations, equipment failures, fluctuating ore quality, and weather, without frequent re-optimization.
+RL allows continuous refinement of dispatching policies to maximize efficiency and resilience in complex, uncertain environments.
+However, RL’s reliance on environmental interaction presents challenges in real mining due to safety, cost, and timeline concerns.
+Unlike controlled simulations real-time training is impractical as RL requires extensive exploration.
+Despite advancements in sample efficiency [17, 18], mining operations still require hundreds to thousands of episodes, hindering real-world deployment.
+Simulators are preferred for safe, cost-effective RL training.
+They enable rapid learning in diverse virtual scenarios, crucial for robust policy development.
+Common RL simulators like OpenAI Gym [19], although simplified, form the basis for algorithm development.
+Simulators are vital in mining optimization due to industry complexity and variability.
+High-fidelity simulations improve accuracy, ensuring learned strategies transfer effectively to real-world operations, enhancing efficiency and safety, making RL a viable tool for mining logistics.
+2.1.4 RL Applications in Truck Dispatch Optimization
+A limited number of studies have applied RL to truck dispatch optimization, but these are typically tailored to specific RL algorithms, limiting their adaptability.
+Adapting them to work with different RL methods would require extensive understanding and potential modifications.
+Huo et al.
+[20] apply Q-learning to optimize dispatching in haulage operations, reducing greenhouse gas emissions while maintaining production.
+Matsui et al.
+[14] develop a real-time dispatching algorithm using deep RL for autonomous haulage trucks, improving transportation efficiency and fuel consumption.
+De et al.
+[21] extend RL to short-term production planning, integrating actor-critic agents for equipment allocation and production scheduling.
+Chiarot et al.
+[15] apply Q-learning-based deep RL to reduce delays during shifts and breaks, improving material supply to crushers.
+Ref. Year Uncertainties Simulator arch. RL adaptability RT Visualization Customizable Platform used OS*
 [22] 2022 Extensive DES N.A. $\times$ $\checkmark$ Simpy $\times$
@@
 Ours 2025 Extensive DES Adaptable Extensive $\checkmark$ Salabim, Python $\checkmark$
-TABLE I: A study of conventional simulators used in Surface Mining process optimization, especially for Truck Dispatching.
-Comparative analysis of mining simulators based on key features.
+Table 1: Comparative analysis of mining simulators for truck dispatching based on key features.
 The table evaluates simulators across several dimensions: Uncertainties, ranging from Extensive (comprehensive modeling of equipment failures, maintenance, variable haul times) to Fair (modeling of limited uncertainties); Simulator Architecture, comparing (DES) and Simpler Rule-based Logic; RL Adaptability, which indicates whether the simulator is Adaptable (native integration), Custom (requires adaptation), or N.A.
 (not designed for RL); Real-time Visualization, measuring the extent of Extensive (Elaborate animated visuals with KPI dashboards), Fair (basic visual representation, or None.
-However, due to the limitations of static optimization methods in dynamic environments, research is increasingly focusing on alternative, adaptive approaches.
-II-A2 DES and integration with RL in Mining Operations
-To effectively apply RL in mining logistics, it must be integrated with robust simulation frameworks.
-DES has been widely used in industrial and mining applications for optimizing processes like truck scheduling.
-By modeling stochastic interactions between equipment and processors, DES captures variability and complexity, often replacing intricate mathematical models with probabilistic parameters [14].
-Applications include supply chain evaluation [15] and ensuring adherence to production schedules [16], demonstrating DES’s role in handling uncertainty and improving operational efficiency.
-Recent research has explored combining RL with DES to simulate complex industrial environments.
-While standard RL environments like OpenAI Gym provide useful benchmarks, they often lack the realism required for industrial applications [17].
-Industrial settings demand detailed, stochastic modeling, similar to DES, to account for dynamic conditions and resource constraints.
-Integrating RL’s trial-and-error learning with DES has gained traction for modeling real-world stochastic systems.
-For example, [18] transformed DES-based SCT controllers into RL environments, enhancing decision-making in automotive plant control.
-This RL-DES integration has been applied to scheduling and optimization across industries.
-[19] employed Deep Q-Networks (DQN) for flexible job shop problems, outperforming traditional metaheuristics, while [20] used DES and OpenAI Gym to create RL-compatible production scheduling environments, simplifying RL algorithm deployment.
-In mining, RL has been explored for short-term planning, truck dispatching, and scheduling.
-[6] introduced a curriculum-driven RL method for vehicle dispatching to address sparse rewards, while [8] developed a real-time RL-based dispatching system for autonomous trucks.
-Additionally, [13] applied Q-learning to optimize material supply during operational delays.
-These studies highlight the potential of RL-DES integration in enhancing decision-making and efficiency in mining operations.
-II-A3 RL and the importance of simulator
-RL excels in sequential decision-making, delivering state-of-the-art performance across various domains, including robotics, locomotion control, autonomous driving, and multi-agent systems.
-[21] The mining truck dispatching problem can be framed as a sequential decision-making task, where truck assignments must adapt to evolving conditions.
-RL provides a suitable framework for optimizing these dispatching strategies
-In mining, RL enables adaptive decision-making by learning optimal dispatching policies through trial-and-error interactions.
-This approach accommodates dynamic changes in configurations, equipment failures, fluctuating ore quality, and weather, without frequent re-optimization.
-RL allows continuous refinement of dispatching policies to maximize efficiency and resilience in complex, uncertain environments.
-However, RL’s reliance on environmental interaction presents challenges in real mining due to safety, cost, and timeline concerns.
-Unlike controlled simulations real-time training is impractical as RL requires extensive exploration.
-Despite advancements in sample efficiency [22, 23], mining operations still require hundreds to thousands of episodes, hindering real-world deployment.
-Simulators are preferred for safe, cost-effective RL training.
-They enable rapid learning in diverse virtual scenarios, crucial for robust policy development.
-Common RL simulators like OpenAI Gym [24], although simplified, form the basis for algorithm development.
-Simulators are vital in mining optimization due to industry complexity and variability.
-High-fidelity simulations improve accuracy, ensuring learned strategies transfer effectively to real-world operations, enhancing efficiency and safety, making RL a viable tool for mining logistics.
-II-A4 RL Applications in Truck Dispatch Optimization
-A limited number of studies have applied RL to truck dispatch optimization, but these are typically tailored to specific RL algorithms, limiting their adaptability.
-Adapting them to work with different RL methods would require extensive understanding and potential modifications.
-Huo et al.
-[7] apply Q-learning to optimize dispatching in haulage operations, reducing greenhouse gas emissions while maintaining production.
-Matsui et al.
-[8] develop a real-time dispatching algorithm using deep RL for autonomous haulage trucks, improving transportation efficiency and fuel consumption.
-De et al.
-[10] extend RL to short-term production planning, integrating actor-critic agents for equipment allocation and production scheduling.
-Chiarot et al.
-[13] apply Q-learning-based deep RL to reduce delays during shifts and breaks, improving material supply to crushers.
-II-A5 Limitations: Existing simulators
-Refer to Table I for a comparison of conventional simulators based on crucial features.
+*OS stands for Open Source/ public availability of code or simulator for use.
+2.1.5 Limitations: Existing simulators
+Refer to Table 1 for a comparison of conventional simulators based on crucial features.
 A notable absence of real-time visualization is observed in several studies, including [22, 24, 13, 20, 14, 3, 21, 15], which can hinder the ability to monitor and manage mining operations effectively.
@@
 Accurate simulation of real-world complexities and uncertainties, along with seamless integration with widely accepted formats like OpenAI Gym, for ease of use and training of off-the-shelf and custom algorithms.
-Figure 1: Overview of the Mining-Gym framework from different perspectives.
-(a) Event-driven or decision-point-based RL setting: The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment.
-(b) Comprehensive system architecture: Displays all key components of the Mining-Gym, including the graphical user interface (GUI), the generated configuration file used to initialize the environment, and the dashboard with real-time visualizations.
-(c) OpenAI Gym-compatible RL interface: Illustrates how the Mining-Gym integrates with OpenAI Gym by adapting simulation signals into the standard reset and step methods.
+Figure 1: (a) Comprehensive system architecture: Displays all key components of the Mining-Gym, including the graphical user interface (GUI), the generated configuration file used to initialize the environment, and the dashboard with real-time visualizations.
+(b) OpenAI Gym-compatible RL interface: Illustrates how the Mining-Gym integrates with OpenAI Gym by adapting simulation signals into the standard reset and step methods.
 This compatibility allows seamless integration with popular RL libraries, such as Stable-Baselines3 [27], enabling easy training and testing of RL models.
-Note: the reward function in (a) is for demonstration only.
-III Mining-Gym: System Architecture and Design
-The mining process especially the dynamics and uncertainty of the Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle is captured using a DES based simulation model.
-While the scheduling problem is modeled as a MDP in order to solve it using RL algorithms (see Fig.1).
-Following we discuss these two modeling processes.
-III-A Mining-Gym overview
-The mining gym simulator and benchmarking tool is a sophisticated platform designed for RL applications in mining operations.
+3 Mining-Gym: System Architecture and Design
+Mining-Gym’s architecture integrates two complementary modeling approaches to tackle the truck dispatch scheduling challenge.
+It captures the dynamic, stochastic Load-Haul-Dump (crusher or dumping site)-Return-Queue (LHDRQ) cycle via a high-fidelity Discrete Event Simulation (DES) model, while the dispatch scheduling is formulated as a Markov Decision Process (MDP), enabling Reinforcement Learning (RL) for adaptive optimization.
+This dual-modeling framework represents both the operational complexity of mining and the sequential decision-making essential for intelligent dispatching.
+The following sections detail Mining-Gym’s architecture, covering the system overview, DES implementation, and MDP formulation for RL-based optimization.
+3.1 Mining-Gym overview
+Mining-Gym is a mining gym simulator and benchmarking tool designed for RL applications in mining operations.
 It features a GUI-based interface where users input crucial parameter values that define the state of the mining site, equipment, and other relevant factors.
-These parameters are then translated into a human-readable and editable text file, allowing for easy adjustments directly within the simulation environment.
+These parameters are translated into a human-readable text file, where they are divided into fixed parameters (sampled once for the entire run, defining the infrastructure) and operational parameters (stochastically sampled per shift/episode, modeling daily variability).
+This structure allows for easy adjustments directly within the simulation environment, while separate scenario files are now used to manage failure conditions for clearer experimentation.
 Users can choose to either train a scheduler policy or run an existing one, which can be either classical or RL-based.
 The simulator includes an additional interface that supports the training and testing of modern RL algorithms, enabling advanced simulation capabilities.
-Furthermore, the tool provides a real-time dashboard displaying key performance indicators (KPIs) and a dynamic, visual representation of the mining site’s operations, offering users valuable insights into the system’s performance and the effects of various parameters on mining activities.
-We present a DES environment for mining processes, integrated with OpenAI Gym—rebranded as Gymnasium in 2023—for RL applications.
+Furthermore, the tool provides a real-time dashboard displaying key performance indicators (KPIs) and a dynamic, visual representation of the mining site’s operations, offering users valuable insights into the system’s performance and the effects of various parameters on mining activities (see Fig.
+1).
+We present a DES environment for mining processes, integrated with OpenAI Gym, (currently Gymnasium) —for RL applications.
 Gymnasium is a widely used toolkit that standardizes interactions across diverse environments, from control tasks to robotics and video games.
@@
 Together, Gymnasium and Stable Baselines create a powerful ecosystem for RL research, facilitating reproducible experimentation and benchmarking in complex domains like mining simulations.
-III-B Mining Process Modeling: DES Overview
-For DES modeling we have used a comprehensive python package named Salabim [26], which offers process interaction methods, queue handling, resources, statistical sampling, and real-time 2D/3D animation capabilities.
+3.2 Mining Process Modeling: DES Overview
+For DES modeling, we have used a comprehensive python package named Salabim  [^2] Salabim is an open-source DES and animation library in Python.
+Available at https://www.salabim.org/ which offers process interaction methods, queue handling, resources, statistical sampling, and real-time 2D/3D animation capabilities.
 Components are fundamental building blocks that define the dynamic behavior of entities within the simulation environment.
@@
 Specifically, in Mining-Gym, Shovels, Dumps and Crushers are modeled as resources.
-Figure 2: (A) Simplified mining system showing three key components: (1) Resource Handler managing resource availability and assignments, (2) Preemption Handler detecting breakdowns and managing repair processes (B) DES-RL interaction flow illustrating how the RL policy integrates with the DES.
-At decision points, the environment state is processed by the RL policy to determine resource assignments.
-Immediate or step rewards guide learning during simulation, while the episodic reward at shift (or episode) end updates the policy before environment reset.
-(C) Load-Haul-Dump-Return-Query (LHDRQ) cycle illustrating the truck’s journey through the mining process, which begins with querying the dispatcher for assignments, followed by loading material, hauling to the destination, dumping, and returning empty.
-Breakdown events, managed by the Preemption Handler, can interrupt operations at any stage.
-In the SALABIM framework, the Truck component In the SALABIM framework, the Truck component follows a Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle.
-In the loading phase, a truck requests an available shovel as advised by the dispatcher and waits its turn.
-Once granted access, it undergoes loading, simulating material transfer time.
+In the Salabim framework, the Truck component follows a Load–Haul– Dump–Return–Query (LHDRQ) cycle with stochastic parameters representing operational variability.
+In the Loading phase, a truck requests an available shovel as advised by the Dispatcher (classical or RL based) and waits for its turn.
+Once granted access, it undergoes loading, with material transfer time sampled from a configurable probability distribution (see Table 8 for example parameterizations).
 The truck then enters the Haul phase, traveling a predefined trajectory to the dump (crusher or dumping site), represented by holding for travel time.
-The dump choice is made at the loading site or shovel.
+Travel time depends on load state: loaded trucks travel at reduced speed due to fuel penalties, while empty returns occur at higher speed.
+Both speeds are sampled from configurable distributions to reflect operational variability.
+At the dump, the truck enters the Dump phase, requesting access to an available crusher or dumping site.
+After access, it performs dumping, with unloading time sampled from distributions that differ between crushers and dumping sites to reflect their operational characteristics.
+The Return phase follows, where the truck travels back to the shovel, completing one haul cycle.
+Stochastic disruptions can interrupt any phase: equipment failures occur according to Mean Time Between Failure (MTBF) distributions (typically modeled as Poisson processes), triggering the preemption handler that suspends operations until repairs complete after a Mean Time To Repair (MTTR) sampled from appropriate distributions.
+The Query phase then begins, where the truck seeks scheduling from the dispatcher.
+This event-driven approach ensures dispatch decisions account for truck capacity constraints, variable cycle times, and disruption-induced uncertainty.
+All distributional parameters are fully configurable through the Mining-Gym interface, allowing users to calibrate the simulation to specific mine characteristics or test algorithmic robustness across diverse operational settings.
 At the dump, the truck enters the Dump phase, requesting access to an available crusher or dumping site.
@@
 Random) aswell as learned strategies (RL based).
-Only one dispatcher can be replaced with a learnable NN-based learnable policy at a time, with shovel allocation set as the default for all experiments.
-Modeling breakdown events is crucial for simulating disruptions when trucks, shovels, or dumps become unavailable due to failures.
-These breakdowns impact loading, dumping, and transit, affecting efficiency and resource use.
-A preemption handler enhances realism by managing resource interruptions.
-For example, if a shovel breaks down mid-loading, the preemption handler pauses operations until repairs or replacement occur, accurately reflecting real-world maintenance disruptions.
-Additional features : in the final Mining-Gym simulation build upon the basic workflow (Fig. 2).
+Figure 2: (A) Simplified mine-site simulation logic of Mining-Gym, showing three key components: (1) Resource Handler managing resource availability and assignments, (2) Preemption Handler detecting breakdowns and managing repair processes (B) Load-Haul-Dump-Return-Query (LHDRQ) cycle illustrating the truck’s journey through the mining process, which begins with querying the dispatcher for assignments, followed by loading material, hauling to the destination, dumping, and returning empty.
+Breakdown events, managed by the Preemption Handler, can interrupt operations at any stage.
+(C) DES-RL interaction flow illustrating how the RL policy integrates with the DES.
+At decision points, the environment state is processed by the RL policy to determine resource assignments.
+Immediate or step rewards guide learning during simulation, while the episodic reward at shift (or episode) end updates the policy before environment reset.
+Algorithm 1 MineGym: Event-Driven DES-RL Synchronization
+1:Simulation shift duration $T_{\text{shift}}$
+2:Trajectories $\{(s_{t},a_{t},r_{t})\}_{t=0}^{T_{\text{shift}}}$
+3:Initialize DES $\mathcal{S}_{\text{DES}}\leftarrow\texttt{Salabim.Environment()}$
+4:Initialize RL Agent $\mathcal{A}_{\text{RL}}\leftarrow\texttt{Gym.Agent()}$
+[…]
+6:while $t_{\text{current}}<T_{\text{shift}}$ do
+7:  // Decision Point Trigger (DES Pauses)
+8:  $\texttt{event}\leftarrow\mathcal{S}_{\text{DES}}.\texttt{run\_until\_next\_decision\_event()}$
+9:  $\tau\leftarrow\mathcal{S}_{\text{DES}}.\texttt{event\_time\_advance}$
+10:  $d\leftarrow\texttt{event.\text{triggering\_entity}}$ (e.g., Truck requesting assignment)
+11:  $\triangleright$ The DES execution is instantaneously paused at $t=\mathcal{S}_{\text{DES}}.\texttt{now()}$
+12:  if $d.\texttt{is\_initial\_state}$ then
+13:    $a_{t}\leftarrow\pi_{\text{classic}}(\text{shovel states})$ $\triangleright$ Use default policy for initialization
+14:  else
+15:    // RL Observation
+16:    $s_{t}\leftarrow\mathcal{S}_{\text{DES}}.\texttt{get\_instantaneous\_state}(d)$ $\triangleright$ Snapshot of system state at exact moment of pause
+17:    // Synchronous Communication (Gym Step)
+18:    $\left(a_{t},r_{t}\right)\leftarrow\mathcal{A}_{\text{RL}}.\texttt{step\_and\_block}(s_{t},\tau)$ $\triangleright$ Agent computes action $a_{t}$, DES thread blocks until action received
+19:    $\triangleright$ Reward $r_{t}$ is computed by the Agent, often using $\tau$
+20:  end if
+21:  // DES Resumes
+22:  $d.\texttt{assign\_resource}(a_{t})$ $\triangleright$ Apply the chosen action to the truck entity
+23:  $\mathcal{S}_{\text{DES}}.\texttt{resume\_and\_advance\_time}(\tau)$
+24:  $t_{\text{current}}\leftarrow t_{\text{current}}+\tau$
+25:end while
+Additional features: The Mining-Gym workflow integrates DES modeling with RL-based decision-making through several key mechanisms illustrated in Fig. 2 and Algorithm 1.
+Core System Architecture (Fig.
+2(A)): The simulation operates through two critical handlers: (1) the Resource Handler manages resource availability, queuing, and assignments for shovels, crushers, and dumps; (2) the Preemption Handler detects equipment breakdowns and coordinates repair processes, ensuring realistic disruption modeling across all equipment types.
+Truck Operation Cycle (Fig.
+2(B)): Trucks follow the Load-Haul-Dump-Return- Query (LHDRQ) cycle, where each phase—from loading at shovels through hauling to dumps/crushers and returning—can be interrupted by breakdown events managed by the Preemption Handler.
+This event-driven architecture captures operational uncertainties inherent in mining logistics.
+DES-RL Synchronization (Fig.
+2(A) & Algorithm 1): At decision points (when trucks query for assignments), the DES pauses and transfers the current system state to the RL policy (Algorithm 1, lines 6-14).
+The policy computes resource assignments while the DES thread blocks (line 16), ensuring synchronous communication.
+Immediate rewards guide step-wise learning, while episodic rewards at shift end enable policy updates before environment reset (lines 17, 22).
+The complete workflow integrates several advanced features designed to enhance realism and flexibility, which are presented as follows.
+  • 1.
     Choice between Dump Types: The stripping ratio of waste to ore is represented by the parameter $\epsilon$, which controls the probability that a truck carrying ore is directed to the crusher rather than the dumping site.
@@
 This parameter can be adjusted to balance waste and ore management, adapting to varying operational conditions.
+  • 2.
     Scalable Configurations: The simulation supports adjustable configurations for Trucks, Shovels, Crushers, and Dumping sites, allowing for scalable operational simulations that can be tailored to different mining scenarios.
+  • 3.
     Uncertainty Modeling: To reflect the stochastic nature of real mining operations, the simulation incorporates sampling from various probability distributions.
 These distributions, along with their parameters, are configurable, enabling flexible modeling of operational uncertainty.
-  • Dispatcher Strategies: Non-learnable, baseline dispatcher strategies (default random) for resources like Crushers and Dumping sites are included.
-These strategies allow for the evaluation of different dispatching approaches without relying on RL.
-III-C Modeling the Dynamic Dispatching Problem as MDP
-III-C1 Typical RL
+  • 4.
+    Dispatcher Strategies: Non-learnable baseline dispatching strategies, including the default random policy, are incorporated.
+These baselines enable systematic evaluation of different dispatching approaches and provide a reference for comparing the performance of RL-based policies
+For detailed modeling assumptions underlying our DES implementation, please refer to Appendix A.
+3.3 Modeling the Dynamic Dispatching Problem as MDP
+3.3.1 Typical RL
 In a typical RL setting, the agent interacts with the environment at every time step $t$.
@@
 The goal is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative reward $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$, where $\gamma$ is the discount factor.
-III-C2 Decision-point based or Event-driven RL
+3.3.2 Decision-point based or Event-driven RL
 In mining truck scheduling, an event-driven reinforcement learning (RL) approach is adopted, where the agent interacts with the environment only at discrete decision points $d$, rather than at every time step.
@@
 in vehicle routing [28] and supply chain management [29].
-III-C3 Defining Agent-Environment interaction and the MDP
+3.3.3 Defining Agent-Environment interaction and the MDP
 Formulating truck dispatch scheduling in open-pit mining as a Markov Decision Process (MDP) involves defining the components of MDP: states, actions, rewards, and transitions.
@@
 The state of the system at a given time must encode all the features needed for the agent to learn a relationship with the desired objective to be maximized.
-For this task the state of the system is encoded as a vector with the following components
-    $s^{d}=[SA_{d},TA_{d}]$ (1)
-    where,
-    $SA_{d}$ represents shovel-related attributes as an encoded vector at the current decision point $d$.
+For this task, the state of the system is encoded as a vector with the following components:
+    $s^{d}=[SA_{d},TA_{d}],$ (1)
+    where $SA_{d}$ represents shovel-related attributes as an encoded vector at the current decision point $d$.
 It includes Shovel ID, encoded in 3-bit chunks, Queue Length, representing the number of trucks waiting per shovel, and Shovel Status, a binary indicator of shovel availability (online/offline).
-Similarly, $TA_{d}$ captures truck-related attributes, including TruckID, encoded in 5-bit chunks, Trips complete, tracking the number of trips per truck, and Trip Status, a multi-bit representation of the truck’s current state, such as loading, transit, or maintenance (see Table: II).
+Similarly, $TA_{d}$ captures truck-related attributes, including TruckID, encoded in 5-bit chunks, Trips complete, tracking the number of trips per truck, and Trip Status, a multi-bit representation of the truck’s current state, such as loading, transit, or maintenance (see Table 2).
     Category Attribute / Status Description
-    Shovels $(SA_{d})$ Shovel ID 3-bit binary
-    Queue Length Normalized float
-    Shovel Status Binary (1 or 0)
-    Trucks $(TA_{d})$ Truck ID 5-bit binary
-    Trips Complete Normalized float
+    Shovels $(SA_{d})$ Shovel ID 4-bit binary (supports up to 16 shovels)
+    Queue Length Normalized float [0, 1] (bounded by $2\times N_{trucks}$)
+    Shovel Status Binary (1=operational, 0=breakdown)
+    Active Truck $(TA_{d,active})$ Truck ID 6-bit binary (supports up to 64 trucks)
+    Trips Complete Normalized float (scaled by 500)
     Truck Status 3-bit binary
-    Truck Status Codes
+    Fleet Context $(FC_{d})$ Fleet Avg Trips Normalized float [0, 1] (mean trips across all trucks)
+    Recent Shovel Usage Vector of floats (proportion of last 10 decisions per shovel)
+    Fleet Diversity Normalized float [0, 1] (Shannon entropy of decisions)
     Truck Status Binary Code
@@
     Breakdown 111
-    TABLE II: State Space Description and Truck Status Codes
+    Table 2: Definition of the state-space representation used in the mining fleet simulation.
+Table (a) outlines the structured attributes describing the current system state, including shovel-specific data $(SA_{d})$, active truck attributes $(TA_{d,active})$, and fleet-level contextual variables $(FC_{d})$.
+Table (b) provides the corresponding 3-bit binary encoding scheme for truck operational statuses, used to represent each truck’s current activity or condition within the model.
   • 2)
@@
     In our framework, an action corresponds to resource or shovel allocation and the action space is defined as:
-    $A_{d}\in\{1,2,\dots,SH\}$
+    $A_{d}\in\{1,2,\dots,SH\},$
     where $SH$ represents the total number of shovels, and the action value indicates the shovel ID to which a truck is assigned at decision point $d$.
@@
     Once the dispatching agent $\mathcal{D}$ outputs action $A_{i}^{d}$, the DES environment responds with:
-    $R_{i}^{d}=r_{\text{imm}}^{d}+\big{(}r_{\text{epi}}^{d}\;\text{if}\;d=d_{T}\;\text{else}\;0\big{)}$ (2)
+    $R_{i}^{d}=r_{\text{imm}}^{d}+\big(r_{\text{epi}}^{d}\;\text{if}\;d=d_{T}\;\text{else}\;0\big),$ (2)
     where $r_{\text{epi}}^{d}$ is the episodic reward at the end of an episode (or shift), and $r_{\text{imm}}^{d}$ is the immediate reward per decision-step $d$.
     Immediate Reward Formulation: The immediate reward uses an exponentially weighted sliding window over the latest $k$ decision points:
-    $r_{\text{imm}}^{d}$ $=-\alpha\cdot\hat{TT}_{\text{Avg}}-\beta\cdot\hat{Q}_{\text{Avg}_{d}}-\gamma\cdot(1-D_{\text{DivScr}})$ (3)
-    $=-\alpha\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot\tau_{j}}{\sum_{j=i-k+1}^{i}w_{j}}-\beta\cdot\frac{\sum_{j=i-k+1}^{i}w_{j}\cdot Q_{{SH}_{j}}}{\sum_{j=i-k+1}^{i}w_{j}}$
-    $\quad-\gamma\cdot(1-D_{\text{DivScr}})$ (4)
-    where $w_{j}=e^{\lambda(j-i+k)}$ represents exponential weighting, and $\alpha,\beta,\gamma$ are weighting parameters such that $\alpha+\beta+\gamma=1$.
-    The components are:
-    Average Trip Time:
-    $\tau_{d}=\frac{1}{|\Omega(d)|}\sum_{\omega\in\Omega(d)}\mathcal{T}_{\omega}(d)$ (5)
-    Shovel Queue Time:
-    $Q_{{SH}_{d}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\frac{1}{|\Omega_{\gamma}(d)|}\sum_{\omega\in\Omega_{\gamma}(d)}Q_{\omega}(d)$ (6)
-    Diversity Score:
-    $D_{\text{DivScr}}=\frac{|\{\gamma\in\Gamma\mid A_{\gamma}>0\}|}{|\Gamma|}$ (7)
-    where $A_{\gamma}$ is the number of trucks assigned to shovel $\gamma$ in the recent window.
-    Episodic Reward Formulation: The episodic reward balances production efficiency and resource utilization:
-    $r_{\text{epi}}^{d_{T}}$ $=\omega_{1}\cdot P_{\text{Efficiency}}-\omega_{2}\cdot D_{\text{Imbalance}}$ (8)
-    $=\omega_{1}\cdot\min\left(1,\frac{P_{\text{Vol}}}{P_{\text{Vol}_{\text{Target}}}}\right)-\omega_{2}\cdot\frac{D_{\text{Penal}}}{\sum_{\gamma\in\Gamma}A_{\gamma}}$ (9)
-    where: $P_{\text{Vol}}$ is total material produced per shift, $P_{\text{Vol}_{\text{Target}}}$ is the production target per shift.
-$D_{\text{Penal}}=\max_{\gamma\in\Gamma}A_{\gamma}-\min_{\gamma\in\Gamma}A_{\gamma}$ measures shovel utilization imbalance, $\omega_{1},\omega_{2}$ are weighting parameters where $\omega_{1}+\omega_{2}=1$.
-    This structure balances short-term efficiency with long-term production goals, guiding RL agents in open-pit mining operations.
-During each episode (e.g., shifts of 12 hours), the agent takes $N_{\text{steps}}$ actions, and the discounted sum of rewards defines the return $G_{t}$:
-    $G_{t}=\sum_{k=t+1}^{N_{\text{steps}}}\gamma^{k-t-1}R_{k}$
-    where $\gamma$ is the discounting factor that determines the impact of future actions on the objective function.
-The objective is to maximize the expected return $G_{t}$ by training the agent to improve actions, ensuring trucks meet production targets and minimize queue formation:
-    $\max_{\pi}\mathbb{E}[G_{t}|S_{t}=s_{t},A_{t}=a_{t}]$
-    Policy: The policy defines the strategy that the agent follows to select actions (shovel IDs) based on the current state of the system.
-Given the state $s_{d}=[SA_{d},TA_{d}]$, which encodes the current shovel and truck attributes, the policy predicts the shovel ID to which a truck should be assigned at decision point $d$.
-The policy maximizes cumulative reward by optimizing truck-shovel assignments for operational efficiency, adapting over time based on rewards to meet short- and long-term production goals, ensuring efficient dispatch and shovel utilization.
-The policy $\pi_{\theta}(a|s)$ is a function that gives the probability of taking action $a$ (i.e., assigning a specific shovel ID) given the current state $s$:
-    $\pi_{\theta}(a|s)=\mathbb{P}(A_{d}=a|S_{d}=s)$
-    where $A_{d}$ is the action taken at decision point $d$, which is the shovel ID assigned to a truck, and $S_{d}=s$ is the state of the system at decision point $d$, including both shovel-related attributes $S_{A_{d}}$ and truck-related attributes $T_{A_{d}}$.
-This ensures that the agent’s actions evolve to maximize long-term efficiency and achieve the desired operational goals.
-IV Implementation Features
-In this section we introduce different visualizations and other GUI-based tools that we provide with the MINING-GYM simulator, for better performance comparison and studying change in KPIs
-IV-A Configuration System
+    3.4 Reward Function Formulation
+    The immediate reward at decision point $d$ incorporates four penalty components:
+    $r^{d}_{\text{imm}}=-\alpha\cdot\overline{TT}_{\text{Avg}}^{\text{norm}}-\beta\cdot\overline{Q}_{\text{Avg}}^{\text{norm}}-\gamma\cdot(1-D_{\text{DivScr}})-\delta\cdot S_{\text{streak}}$ (3)
+    where $\alpha,\beta,\gamma,\delta$ are component weights.
+Min-max normalization is applied to the exponentially-weighted averages: $x^{\text{norm}}=\frac{x-x_{\text{min}}}{x_{\text{max}}-x_{\text{min}}}$
+    $\overline{TT}_{\text{Avg}}=\sum_{j=1}^{k}\bar{w}_{j}\cdot\tau_{j},\quad\overline{Q}_{\text{Avg}}=\sum_{j=1}^{k}\bar{w}_{j}\cdot Q^{SH}_{j},\quad\bar{w}_{j}=\frac{e^{\mu\cdot j}}{\sum_{i=1}^{k}e^{\mu\cdot i}}$ (4)
+    where $n=\min(k,|D|)$ is the effective window size with decay parameter $\lambda$. The components are:
+    • (a)
+      Average Trip Time: $\tau_{d}=\frac{1}{|\Omega(d)|}\sum_{\omega\in\Omega(d)}\mathcal{T}_{\omega}(d)$, where $\Omega(d)$ is the set of trucks completing trips and $\mathcal{T}_{\omega}(d)$ is the cycle time for truck $\omega$.
+    • (b)
+      Shovel Queue Time: $Q_{{SH}_{d}}=\frac{1}{|\Gamma|}\sum_{\gamma\in\Gamma}\frac{1}{|\Omega_{\gamma}(d)|}\sum_{\omega\in\Omega_{\gamma}(d)}Q_{\omega}(d)$, where $\Gamma$ is the set of shovels and $Q_{\omega}(d)$ is the waiting time for truck $\omega$ at shovel $\gamma$.
+    • (c)
+      Diversity Score: $D_{\text{DivScr}}=H/H_{\max}$, where $H=-\sum_{i=1}^{|\Gamma|}p_{i}\log_{2}(p_{i}+\epsilon)$, $H_{\max}=\log_{2}(|\Gamma|)$, and $p_{i}=\text{count}_{i}/\sum_{j}\text{count}_{j}$ represents the proportion of assignments to shovel $i$ within a sliding window of size $|W_{\text{div}}|$.
+This Shannon entropy-based measure quantifies shovel utilization uniformity.
+    • (d)
+      Streak Penalty: $S_{\text{streak}}=\frac{1}{|W_{\text{streak}}|-1}\sum_{i=1}^{|W_{\text{streak}}|-1}\mathbb{1}(a_{i}=a_{i-1})$, where $W_{\text{streak}}$ represents the recent $|W_{\text{streak}}|$ decisions, $a_{i}$ is the assigned shovel at decision $i$, and $\mathbb{1}(\cdot)$ is the indicator function.
+This penalizes consecutive assignments to the same shovel.
+    The reward function parameters used in the demonstration experiments were determined through preliminary tuning to balance multiple operational objectives: production efficiency, resource utilization uniformity, and queue management.
+Episodic production weight $\omega_{1}=0.4$ and diversity penalty weight $\omega_{2}=0.6$ were applied.
+A high–performance bonus $b_{1}=0.5$ was triggered when $P_{\text{ratio}}\geq 0.95$ and $D_{\text{final}}>0.65$, while a moderate–performance bonus $b_{2}=0.2$ was triggered when $P_{\text{ratio}}\geq 0.90$ and $D_{\text{final}}>0.50$.
+Immediate rewards used a $k=5$ decision–point sliding window with exponential decay $\mu=0.5$.
+    Complete parameter specifications are documented in Table B.9 (Appendix B).
+These hyperparameter values are provided as baseline configurations; Mining-Gym users can readily adjust these parameters through configuration files to reflect domain-specific operational priorities or conduct ablation studies on individual reward components.
+    The episodic reward at shift termination evaluates overall performance:
+    $r_{\text{epi}}^{d_{T}}=\omega_{1}\cdot P_{\text{ratio}}-\omega_{2}\cdot(1-D_{\text{final}})+B_{\text{perf}}$ (5)
+    where $P_{\text{ratio}}=\min(1,P_{\text{Vol}}/P_{\text{Vol,Target}})$ is the production ratio, $D_{\text{final}}$ is the final diversity score, and the performance bonus is:
+    $B_{\text{perf}}=\begin{cases}b_{1}&\text{if }P_{\text{ratio}}\geq\theta_{1}\text{ and }D_{\text{final}}>\phi_{1}\\ b_{2}&\text{if }P_{\text{ratio}}\geq\theta_{2}\text{ and }D_{\text{final}}>\phi_{2}\\ 0&\text{otherwise}\end{cases}$ (6)
+    with bonus values $b_{1}>b_{2}>0$ and thresholds $\theta_{1}>\theta_{2}$, $\phi_{1}>\phi_{2}$. The total episode return is:
+    $G=\sum_{d=1}^{d_{T}-1}\gamma^{d-1}r_{\text{imm}}^{d}+\gamma^{d_{T}-1}\left(r_{\text{imm}}^{d_{T}}+r_{\text{epi}}^{d_{T}}\right)$ (7)
+    where $\gamma\in[0,1]$ is the discount factor, and the RL objective maximizes expected return: $\max_{\theta}\mathbb{E}_{\tau\sim\pi_{\theta}}[G]$.
+4 Implementation Features
+In this section, we introduce different visualizations and other GUI-based tools that we provide with the Mining-Gym simulator, for better performance comparison and studying change in KPIs.
+4.1 Configuration System
 Variables Source
@@
 At Shovel, Crusher, Dumping site $Q_{SH},Q_{CR},Q_{DS}$ DES
-Load per trip $L$
+Load per trip $L$ -
 Truck carrying Waste, Ore $L_{W},L_{O}$ User
@@
 Others: -
-Number of trips (SH-CR-DS-SH) $N$ DES
+Number of trips (SH-CR-DS-SH) NN DES
 Available (operation unit) time $T_{\text{SHF}}$ User
@@
 Truck with Waste, Ore, Empty $F_{W},F_{O},F_{E}$ User
-Equipment off time $OF_{E}$
+Equipment off time $OF_{E}$ -
 Truck, Shovel, Crusher, Dumping Site $OF_{\text{TR}},OF_{\text{SH}},$ DES
 $OF_{\text{CR}},OF_{\text{DS}}$
-Equipment idle time $IDL_{E}$
+Equipment idle time $IDL_{E}$ -
 Truck, Shovel, Crusher, Dumping Site $IDL_{\text{TR}},IDL_{\text{SH}},$ DES
@@
 Target Production volume $P_{\text{Vol, Targ}}$ User
-TABLE III: Overall list of variables and parameters used
-Figure 3: (a) Shows a sample page from the GUI interface and the resulting configuration file (b) Shows the Real-time representative visualization of the mine-site.
-The screenshot displays trucks queued at shovels, with the rightmost shovel offline.
-It also shows trucks moving between the dumping site, crushers, and the shovels.
-As discussed in the previous section, a large number of parameters must be initialized to run the simulator (see Table III and Table V).
+Table 3: List of variables and parameters used. See Appendix B for detailed parametrization and distributions.
+Figure 3: (a) Mining-Gym graphical user interface (GUI) and the generated configuration file.
+(b) Real-time representative visualization of the mine site.
+The screengrab shows trucks queued at shovels, with the rightmost shovel offline, and trucks transiting between dump sites, crushers, and shovels.
+As discussed in the previous section, a large number of parameters must be initialized to run the simulator (see Table 3).
 To provide an organized method for supplying this information, we have designed a GUI interface.
@@
 3(a)).
-IV-B Other Visualization Features
+This file now segregates parameters into fixed infrastructural properties (sampled once per experiment, e.g., number of shovels ) and operational variabilities (stochastically sampled per shift, e.g., fuel consumption and MTBF ).
+Furthermore, to ensure standardized and reproducible testing across diverse challenges, failure conditions are managed through a dedicated scenario configuration file.
+This file defines a set of pre-defined stress tests (e.g., loss of one shovel and fifteen trucks) where specific failure parameters are set precisely, allowing for systematic and comparable evaluation of scheduler robustness.
+4.2 Other Visualization Features
 We provide a real-time representative visualization of the mining setup as configured in the setup file.
@@
 3(b)).
-Key Performance Indicators (KPIs) in mining process optimization include metrics like equipment utilization, cycle time, production throughput, fuel efficiency, and downtime.
-These metrics help evaluate the efficiency, productivity, and cost-effectiveness of mining operations, guiding improvements and strategic decisions to enhance overall performance and profitability.
-A key factor in KPI measurement is the time frame, which can range from shifts to daily, weekly, monthly, or yearly intervals.
-By default, we consider a 12-hour shift as the standard, but this can be customized as a parameter.
-We include the following essential KPIs :
+Key Performance Indicators (KPIs): We measure mining efficiency through these essential metrics:
   1. 1.
-    Total Production ($P_{VOL}$): Refers to the total quantity of material (ore + waste) excavated per shift. It is measured in tons, indicating the scale and efficiency of production activities.
-    $P_{VOL}$ $=N\times L$
-    where: $N$ = Total number of trips in a shift, L = Material load per trip (tons)
+    Total production $(\text{shift}^{-1})$: $P_{Vol}=N\times LO$, where $N$ is the total trips completed and $LO$ is ore load per trip (tons).
   2. 2.
-    Trips per Hour ($\text{TPH}_{t}$): The “Trips per Hour” metric quantifies the rate at which the full haul cycle (SH-CR-DS-SH) is completed, normalized to an hourly basis.
-It provides a standardized measure of fleet productivity by capturing the number of new trips executed within the most recent hourly interval.
-    $\text{TPH}(t)=\frac{\Delta t}{60}\left[N(t)-N(t-\Delta t)\right]$
-    where $N(t)$ = Total number of trips (SH-CR-DS-SH) completed by all trucks up to time $t$, $N(t-\Delta t)$ = Total trips at $t-\Delta t$, $\Delta t$ = Measurement interval in minutes (typically 60).
-    Total Number of Trips:
-    $N(t)=\sum_{i=1}^{TR}C_{i}(t)$
-    where $TR$ = Total number of trucks in the fleet, $C_{i}(t)$ = Trips completed by truck $i$ up to time $t$.
+    Trips $(\text{hour}^{-1})$ : $\mathcal{T}_{h}=\frac{60}{\Delta t}\times[N(t)-N(t-\Delta t)]$, where $N(t)$ = cumulative trips at time $t$, $N(t-\Delta t)$ = trips at time $t-\Delta t$, $\Delta t$ = measurement interval (minutes), $h$ is the hour index and $N(t)=\sum_{i=1}^{TR}C_{i}(t)$ with $TR$ = total trucks, $C_{i}(t)$ = trips by truck $i$
+  3. 3.
+    Max.
+queue length $(\text{hour}^{-1})$ : $\widehat{Q}_{SH,h}=\max_{k=1}^{K}\left(\max_{i=1}^{N_{SH}}Q_{SH,i}^{(k)}\right)$, where $h$ is a given hour, $N_{SH}$ the number of shovels, $\mathbf{Q}_{SH}^{(k)}$ the shovel-queue vector at time instant $k$ with element $Q_{SH,i}^{(k)}$ for shovel $i$ ($i=1,\dots,N_{SH}$), and $K$ the number of time instants in hour $h$.
+Measures bottleneck severity: a high value indicates major congestion events (i.e., long queues) occurred, which points to a potential failure point or poor real-time dispatching.
+  4. 4.
+    Avg.
+queue length $(\text{hour}^{-1})$ : $\overline{Q}_{SH,h}=\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{N_{SH}}\sum_{i=1}^{N_{SH}}Q_{SH,i}^{(k)}\right)$.
+Let $h$ be a given hour, $N_{SH}$ the number of shovels, $\mathbf{Q}_{SH}^{(k)}$ the shovel-queue vector at time instance $k$ with element $Q_{SH,i}^{(k)}$ for shovel $i$ ($i=1,\dots,N_{SH}$), and $K$ the number of time-instance in hour $h$.
+This metric measures system utilization: a low value suggests trucks and shovels are well-utilized, while a high value suggests overall system congestion.
+  5. 5.
+    Cost per ton $(\text{shift}^{-1})$ : $\text{CPT}=\dfrac{C_{KW}+C_{EST}}{P_{Vol}}$
+    where $C_{KW}$ = known costs, $C_{EST}$ = estimated costs
+  6. 6.
+    Fuel consumption : $FC=\frac{N\times F_{\text{Unit}}}{P_{\text{Vol}}}$
+    where $F_{\text{Unit}}=F_{W}+F_{O}+F_{E}$ with $F_{W}$, $F_{O}$, $F_{E}$ = fuel for waste, ore, empty trips respectively
 Figure 4: Conceptual Diagram of Minesite
+5 Experimental Setup
+5.0.1 Environment Design
+To address scalability and computational efficiency, we configure a large-scale operational scenario with 8 shovels, 60 trucks, 5 crushers, and 3 dumping sites.
+This represents an upper–medium to large bauxite operation (approximately 5–10 million tonnes annual capacity).
+The resulting truck–to–shovel ratio of 7.5:1 is within typical industry practice (6–12 trucks per shovel), generating sufficient event density for stress-testing.
+With 60 trucks completing 4–6 trips per shift at 100 t payloads, this setup supports approximately 24,000–36,000 t of daily material movement across three shifts.
+More importantly for benchmarking, this scale produces high event density: concurrent truck movements, simultaneous loading/dumping, and overlapping breakdown–repair cycles stress-test both the DES engine and RL decision-making under realistic operational loads.
+This configuration also enables controlled scalability studies across: (i) computational performance under dense event streams, (ii) algorithmic robustness under resource degradation (e.g., losing 15–30 trucks or 2–3 shovels), and (iii) queue-management complexity under realistic contention.
+The scale selected represents an upper bound where RL training remains tractable (5̄00 episodes within reasonable runtimes) while still capturing the operational characteristics and challenges of industrial-scale mining logistics.
+Equipment specifications are listed in Table 4, and a conceptual diagram of the minesite is presented in Fig. 4.
+Equipment Model/ Make Quantity Specification (each)
+Shovels CAT 6040 8 43.7 t bucket size
+Trucks CAT 777 60 98.2 nominal payload
+Crushers Metso Lokotrack (LT106) 5 450 t / hr capacity
+Dumping site - 2 Unlimited
+Table 4: Equipment List with Specifications
+5.0.2 Training Setup
+To ensure that the trained agent generalizes well across the testing scenarios, we not only incorporate stochastic parameters to randomize breakdown and repair times for mining equipment and trucks but also enhance the reward function by introducing queue buildup metrics (queue length) and diversity score to prevent repeatedly selecting same shovels.
+For training an RL policy, we have used the PPO [30] algorithm from Stable Baselines3 [27]).
+We have used all the default parameters and training settings as in the repository’s PPO implementation.
+Note the scenario parameters which help specify the failure scenarios to be simulated during training or testing.
+To validate Mining-Gym as a benchmarking platform and highlight the potential of RL-based truck dispatch scheduling, we present comparative results using a PPO policy trained for 500 episodes (shifts).
+While a proof-of-concept with limited training, these results suffice to demonstrate both the framework’s feasibility and the performance potential of learned scheduling policies versus classical heuristics.
+5.0.3 Testing Setup
+We designed scenarios to test the efficacy of the RL trained scheduler against a vanilla scheduler (random scheduler) as baseline, under different challenging scenarios.
+For the random scheduler, the dispatcher’s action $A_{d}$ is selected randomly from the available set of shovel IDs $\{1,2,\dots,SH\}$. The selection follows a uniform distribution,
+$P(A_{d}=a|S_{d}=s)=\frac{1}{SH}\quad\text{for all }a\in\{1,2,\dots,SH\},$
+i.e. any state $S_{d}=s$, each action $A_{d}=a$ (where $a$ is a shovel ID) is equally likely to be chosen, with a probability of $\frac{1}{SH}$.
+The fixed scheduler employs a deterministic round–robin allocation strategy. Each truck $i$ is permanently assigned to a specific shovel based on its ID using a modulo operation:
+$A_{d,i}=(i-1)\bmod SH$ (8)
+where $i\in\{1,2,\ldots,TR\}$ represents the truck ID, and $SH$ is the total number of shovels. This creates a static, pre–determined mapping:
+$P(A_{d,i}=a\mid S_{d}=s)=\begin{cases}1,&\text{if }a=(i-1)\bmod SH\\ 0,&\text{otherwise}\end{cases}$ (9)
+Unlike the random scheduler, the fixed scheduler’s allocation is independent of the current system state $S_{d}$ and remains constant throughout operation, distributing trucks evenly across shovels in a cyclic pattern.
+In contrast, the RL policy $\pi_{\theta}(a|s)$ is a learned policy, where the action $A_{d}=a$ is selected based on the current state $S_{d}=s$ and the parameters $\theta$. The policy is defined as:
+$\pi_{\theta}(a|s)=P(A_{d}=a|S_{d}=s).$
+This probability is learned through RL algorithm where the agent adjusts $\theta$ based on rewards from its actions, optimizing its decision-making process over time.
+We tested the schedulers under different challenging scenarios, which are governed by a dedicated scenario configuration file to ensure consistent parameter selection and repeatability across all comparative trials.
+Table 5: Testing Scenarios for Scheduler Robustness Evaluation.
+Scenario Failures Operational Objective
+(#SH, #TR)
+A (0, 0) Baseline: All equipment available (No failures).
+B (0, 15) Moderate Truck Stress: 25% of the haul fleet is lost.
+C (1, 0) Single Shovel Loss: Tests response to $\sim$12.5% loading capacity bottleneck.
+D (1, 15) Combined Stress: Simultaneous constraints on both loading and hauling capacity.
+E (2, 30) Severe Disruption: Pushes the system to critical limits (25% SH, 50% TR lost).
+F (3, 0) Critical Bottleneck: Tests scheduling under severe $\sim$37.5% loading capacity loss.
+These scenarios systematically vary the number and type of resource failures, replacing the manual parameter definitions.
+We tested the schedulers under different challenging scenarios, which are governed by a dedicated scenarios to ensure consistent parameter selection and repeatability across all comparative trials.
+These scenarios systematically vary the number and type of resource failures.
+All failure scenarios are repeated according to a over ten repeat trials, with the inherent stochasticity in event timings due to probabilistic distributions.
+These repeats are then considered in the results section for statistical analysis and discussion.
+Complete experimental scenario parameters, including detailed equipment specifications, stochastic parameters, state space descriptions, and technical specifications for trucks, shovels, and crushers can be found in Appendix C.
+Figure 5: Comparison of “Mean Trips per hour” and “Mean Truck Queue Length” over a 6-hour shift under six failure scenarios: (a) No failures $\{0,0\}$, (b) Moderate Truck Stress $\{0,15\}$, (c) Single Shovel Loss $\{1,0\}$, (d) Combined Stress $\{1,15\}$, (e) Severe Disruption $\{2,30\}$, and (f) Critical Bottleneck $\{3,0\}$.
+Shovel failures occurs between 100-150 mins and Truck failures occur between 150-200 mins across the scenarios.
+Scenarios progress from baseline operation to severe capacity constraints on loading and hauling.
+6 Results and Discussion
+[…]
+6.1 Comparing Performance, Efficiency and behavior
+The RL-PPO scheduler demonstrates clear superiority over both fixed and random approaches across all six operational scenarios (A to F), showing the efficacy of RL based dynamic scheduling solution.
+6.1.1 Productivity Performance :
+The RL scheduler achieves an average improvement of $5.7\%$ in total trips completed compared to the combined performance of fixed and random schedulers.
+This gain exhibits remarkable consistency across diverse operational conditions, ranging from $4.1\%$ in Scenario C (single shovel loss) to $8.1\%$ in Scenario D (combined shovel loss with moderate truck stress).
+The most substantial advantages emerge under Combined Stress (Scenario D) and Critical Bottleneck (Scenario F) conditions, demonstrating the RL agent’s ability to maximize throughput through superior real-time assignment decisions when resources are constrained.
+6.1.2 Operational Efficiency :
+The efficiency gains are even more pronounced than productivity improvements.
+The RL approach reduces the average mean maximum queue length by approximately $24.4\%$ across all scenarios.
+The magnitude of these reductions varies with operational constraints.
+The largest improvements occur in scenarios with hauling capacity limitations: Moderate Truck Stress (Scenario B) and Combined Stress (Scenario D), where RL reduces queue lengths by $45.2\%$ and $36.9\%$, respectively.
+These reductions highlight the RL scheduler’s strength in efficiently utilizing limited haul fleet resources and optimally balancing truck distribution.
+Scenarios with hard loading bottlenecks (C, E, F) exhibit more modest reductions, such as $4.4\%$ in Scenario E, as some queuing becomes unavoidable.
+Nevertheless, RL minimizes unnecessary waiting and significantly outperforms the random scheduler.
+These results demonstrate that RL provides a dual benefit: maximizing system output while improving resource efficiency, proving valuable across all operational states from baseline to severe disruptions.
+6.1.3 Behavioral Characteristics of Scheduling Approaches :
+The behavioral patterns of each scheduling method reveal fundamental differences in their operational dynamics.
+The RL scheduler’s superior performance stems from its immediate and effective response to system disturbances.
+At system start (Hour 0, initial queue of 52.0 trucks), RL demonstrates rapid stabilization.
+By Hour 1, RL reduces the average queue to $11.18$, representing a $44.8\%$ improvement compared to Fixed ($16.24$) and Random ($19.85$) schedulers.
+Despite substantial queue range across the six-hour period (reflecting the severity of Scenarios E and F), the RL scheduler consistently maintains the lowest mean queue length ($14.14$), indicating a stable allocation policy that adapts effectively without excessive oscillation.
+The Fixed scheduler shows pronounced instability, evidenced by abrupt oscillations in queue length and a high queue range of $24.89$.
+Under failure scenarios C–F, simple rule logic cannot compensate for dynamic capacity loss, producing rapid queue spikes followed by sharp overcorrections.
+This cyclical behavior highlights a structural limitation of static scheduling in non-stationary systems.
+The Random scheduler performs worst overall, with the highest mean queue length ($22.43$) and large queue range ($18.11$), reflecting the absence of coordinated decision logic.coordination.
+SID Performance ($\text{Mean}\pm\text{SD}$) Statistical Results (Cohen’s d)
+RL Random Fixed Schedule vs. Random vs. Fixed
+A 9000.00 $\pm$ 161.60 8111.11 $\pm$ 518.28 8877.56 $\pm$ 319.29 $2.392$ $0.987$
+B 8622.22 $\pm$ 192.21 7900.00 $\pm$ 614.41 8555.56 $\pm$ 245.52 $1.587$ $0.21$
+C 8433.33 $\pm$ 187.08 8033.11 $\pm$ 557.77 8380.89 $\pm$ 351.58 $1.015$ $0.31$
+D 8566.67 $\pm$ 158.11 7622.22 $\pm$ 834.83 8377.78 $\pm$ 370.06 $1.572$ $0.664$
+E 7744.44 $\pm$ 339.53 7388.89 $\pm$ 527.84 7677.78 $\pm$ 164.15 $0.801$ $0.450$
+F 7822.22 $\pm$ 156.35 7622.22 $\pm$ 523.87 7766.67 $\pm$ 180.28 $0.517$ $0.329$
+Table 6: Mean $\pm$ SD performance of algorithms.
+Cohen’s d effect size is reported for RL compared with Random and Fixed schedules.
+Positive d indicates RL superiority; $|d|<0.2$ indicates comparable performance.
+6.2 Comparing Throughput
+[…]
+7 Conclusion and Future Work
+Mining-Gym addresses a key gap in mining process optimization by offering a configurable, open-source benchmarking environment for evaluating reinforcement learning (RL) algorithms in truck dispatch scheduling.
+It combines high-fidelity discrete-event simulation (DES) modeling, seamless integration with OpenAI Gym, and comprehensive visualization tools.
+Empirical evaluation across six operational scenarios demonstrates that the RL-PPO scheduler achieves an average improvement of 5.7% in total trips completed compared to fixed and random scheduling baselines, with gains ranging from 4.1% under single-asset failure to 8.1% under combined stress conditions.
+The RL approach delivers substantial efficiency improvements, reducing mean maximum queue length by approximately 24.4% on average, with reductions reaching 45.2% .
+These productivity and efficiency gains translate to consistently superior throughput performance, with the RL scheduler demonstrating rapid stabilization and adaptive resource allocation even under severe multi-asset disruptions.
+Mining-Gym promotes reproducible research and supports RL adoption in industrial mining applications.
+Future development will focus on several key areas to enhance Mining-Gym’s capabilities and research utility.
+Enhanced Fidelity will incorporate traffic modeling, terrain features, heterogeneous fleets, and operational constraints to improve simulation realism.
+Scenario Expansion will add test cases that reflect challenges such as weather variability, ore heterogeneity, and dynamic production targets.
+Multi-objective Benchmarks will enable evaluation of trade-offs among production, cost, equipment longevity, and environmental impact through configurable objective weighting and standardized multi-criteria evaluation protocols.
+Robustness Evaluation Framework will implement standardized uncertainty quantification protocols including systematic parameter perturbation testing, Monte Carlo robustness analysis, and statistical confidence interval reporting to support algorithmic research in risk-sensitive decision-making.
+Finally, Digital Twin Integration will connect simulations with real-time operational data for data-driven optimization and practical deployment.
+Acknowledgments
+This research was partly supported by the Advance Queensland Industry Research Fellowship AQIRF024-2021RD4.
+Appendix A Modeling Assumptions and Abstractions
+This section presents the different assumption and abstractions used to simplify the modeling process, in the current version of the simulator
+  1. 1.
+    Traffic setting and routing: The current version of the simulation does not incorporate complex vehicle-traffic scenarios.
+In the future iterations we will aim to integrate this functionality.
+Currently the simulation assumes a bidirectional and perfectly routed traffic network within the DES modeling, without no traffic congestion.
+  2. 2.
+    Geographical Distribution and Equipment Heterogeneity: Shovels are distributed across three geographical zones, with zone-specific travel times to crushers and dumps modeled through location cluster assignments.
+Additionally, shovels have performance classes that determine loading efficiency, affecting cycle times.
+These characteristics are not observable to the RL agent but influence the underlying system dynamics (travel times and loading times), creating a partially observable environment where the agent must learn optimal dispatching without directly knowing individual shovel capabilities or locations.
   3. 3.
-    Cost per Ton ($CPT$): Refers to the financial expenditure incurred for every unit weight (typically one ton) of material produced, processed, or transported.
-    $CPT$ $=\frac{C_{KW}+C_{EST}}{P_{VOL}}$
-    where $C_{KW}$ = Total known cost and $C_{EST}$ = Total estimated cost
+    Trailing transport: Truck cycles transporting tailings to the dumping site are not considered in the current version of the simulator, as they have lower priority and often use separate equipment.
+Including tailings would add unnecessary complexity without significantly improving the efficiency of primary ore transport operations.
   4. 4.
-    Fuel Consumption ($FC$): measures the amount of fuel consumed by mining equipment or vehicles relative to the total material hauled during operations:
-    $FC=\frac{\text{Total fuel consumed}}{P_{\text{Vol}}}$
-    where the Total fuel $\text{Total fuel consumed}=N\times F_{\text{Unit}}$, and the $F_{\text{Unit}}=F_{\text{Unit}}=F_{W}+F_{O}+F_{E}.$
-    Here, $N$ is the number of trips completed by all trucks, $F_{W}$, $F_{O}$, and $F_{E}$ are the fuel consumptions for waste, ore, and empty trips, respectively, and $P_{\text{Vol}}$ is the target production volume.
-We have also provided an interactive dashboard as part of this benchmark to analyze change in these KPIs in real time.
-V Experimental Setup
-V-1 Environment Design
-We consider a medium sized Bauxite ($5-10$ Million Tons output) mine with 3 Shovels, 20 Trucks, 4 Crushers and 2 Dumping sites, detailed specifications are presented in Table: IV and V.
-A conceptual diagram of the minesite is provide in Fig.
-4.
+    Environment Reset: During RL-based training, the environment operates in shifts (episodes).
+Each shift begins with trucks scheduled to shovels using the Equal-Queue scheduler by default.
+The shift ends when the time limit is reached, ignoring trucks still in transit.
+Only completed trips and transported material within the shift’s time frame are counted.
+Appendix B Experimental Scenario Parameters
+Table 7: Equipment List with Specifications
 Equipment Model/Make Quantity Specification (each)
@@
 Dumping site - 2 Unlimited
-TABLE IV: Equipment List with Specifications
-V-2 Training Setup
-To ensure that the trained agent generalizes well across the testing scenarios, we not only incorporate stochastic parameters to randomize breakdown and repair times for mining equipment and trucks but also enhance the reward function by introducing queue buildup metrics (queue length) and diversity score to prevent repeatedly selecting same shovels.
-The parameter values used during training is provided in Table V.
-For training an RL policy we have used the PPO [29]algorithm from StableBaselines3 [25]).
-We have used all the default parameter and training settings as present in the repository’s PPO implementation.
-Note the scenario parameters which help specify the failure scenarios to be simulated during training or testing.
-For this proof-of-concept implementation and demonstration the results of only $400$ episodes or shifts of training.
-Figure 5: Comparison of “Trips per hour” over the shift under different failure scenarios: (a) No failure, (b–d) $(SH_{0},TR_{0})$ values where $SH_{0}$ (Num.
-Shovels failed) and $TR_{0}$ (Num.
-Trucks failed) are set as follows—(b) $(0,6)$, (c) $(1,0)$, (d) $(1,6)$.
-Solid lines represent mean values, and shaded regions indicate variance from repeated runs.
-For shovels, failures initiate between 100–300 minutes ( 2–5 hours), and for trucks, between 100–500 minutes ( 2–8 hours) within a 12-hour shift.
-V-3 Testing Setup
-We designed scenarios to test the efficacy of the RL trained scheduler against a vanilla scheduler (random scheduler) as baseline, under different challenging scenarios.
-For a random scheduler, the dispatcher’s action $A_{d}$ is selected randomly from the available set of shovel IDs $\{1,2,\dots,SH\}$. The selection follows a uniform distribution,
-$P(A_{d}=a|S_{d}=s)=\frac{1}{SH}\quad\text{for all }a\in\{1,2,\dots,SH\}$
-i.e.
-any state $S_{d}=s$, each action $A_{d}=a$ (where $a$ is a shovel ID) is equally likely to be chosen, with a probability of $\frac{1}{SH}$.
-In contrast, the RL policy $\pi_{\theta}(a|s)$ is a learned policy, where the action $A_{d}=a$ is selected based on the current state $S_{d}=s$ and the parameters $\theta$.
-The policy is defined as:
-$\pi_{\theta}(a|s)=P(A_{d}=a|S_{d}=s)$
-This probability is learned through RL algorithm where the agent adjusts $\theta$ based on rewards from its actions, optimizing its decision-making process over time.
-We tested the schedulers under different challenging scenarios:
-  1. 1.
-    All equipment available, represented as $(SH_{0},TR_{0})=(0,0)$, i.e., no Shovels and Trucks offline.
-  2. 2.
-    Six Trucks under maintenance, i.e., $(SH_{0},TR_{0})=(0,6)$.
-  3. 3.
-    One Shovel offline, i.e., $(SH_{0},TR_{0})=(1,0)$.
-  4. 4.
-    One Shovel and Six Trucks unavailable, i.e., $(SH_{0},TR_{0})=(1,6)$.
-All failure scenarios are repeated according to a consistent temporal schedule over five repeat trials, with the inherent stochasticity in event timings due to probabilistic distributions.
-These repeats are then considered in the results section for analysis and discussion.
-Stochastic Parameters Probability Distributions
-Fuel consumption Liters
-Loaded Truck (FL) in lt/ton-km $\text{Normal}(0.25,0.001)$
-Empty Truck (FE) $\text{Normal}(0.35,0.001)$
-Shovel (Lt/hr) $\text{Uniform}(250,300)$
-Equipment specifications Value
-Truck payload (TP) $\text{Normal}(98,0.5)$ t
-Shovel bucket capacity (SCP) $\text{Normal}(43,0.5)$ t
-Empty Truck Speed (TS) $\text{Normal}(40,60)$ km/hr
-Loaded Truck Speed (TL) $\text{Normal}(25,40)$ km/hr
-Equipment Maintenance** Time (in mins )
-MTBF Shovel (FSH) $\text{Poisson}(2520)$
-MTBF Truck (FTR) $\text{Poisson}(2160)$
-MTBF Crusher (FCR) $\text{Poisson}(30000)$
-MTBF Dumping Site (FDS) $\text{Poisson}(60000)$
-MTTR Shovel (RSH) $\text{Poisson}(240)$
-MTTR Truck (RTR) $\text{Poisson}(300)$
-MTTR Crusher (RCR) $\text{Poisson}(240)$
-MTTR Dumping Site (RDS) $\text{Poisson}(120)$
-Loading and Maneuver Time Time (in mins.)
-Time Truck Loading (TRL) $\text{Normal}(8,2)$
-Travel Time Shovel to Dump (STD) $\text{Normal}(15,5)$
-Time Truck Dumping (TRDM) $\text{Normal}(5,1)$
-Travel Time Dump to Shovel (DTS) $\text{Normal}(15,5)$
-Travel Time Shovel to Crusher (STC) $\text{Normal}(15,2)$
-Time Truck Unloading at Crusher (TRCR) $\text{Normal}(5,1)$
-Travel Time Crusher to Shovel (CTS) $\text{Normal}(15,2)$
-Scenario Parameters -
-Shovel To Fail (STF) Binomial (2, 0.5)
-Trucks To Fail (TTF) Binomial (6, 0.5)
-Shovel Initial Breakdown (SIB) Uniform (100,300)
-Truck Initial Breakdown (TIB) Uniform (100,500)
-TABLE V: Distributions for stochastic parameters used in the example. MTBF: Mean Time Between Failures, MTTR: Mean Time To Repair.
-Figure 6: Comparison of production volume $P_{vol}$ in metric tonnes per shift, averaged over 5 repeats trials.
-VI Results and Discussion
-The RL scheduler demonstrated significant performance improvements over the random scheduler across all scenarios, see Fig.5 and Fig. 6 :
-  • Trips Per Hour: RL increased productivity by 18.2% on average, with the greatest improvement (27.2%) observed in the most challenging scenario (1 shovel and 6 trucks offline).
-  • Production Volume: RL delivered 19.7% higher production volumes on average, with improvements ranging from 14.3% to 23.3% across all scenarios.
-The performance advantage of RL was most pronounced in scenarios with multiple resource constraints, suggesting that RL’s optimization capabilities become more valuable as operational complexity increases.
-As is showing through the gain in production volume (see Fig.6) RL improved production volume across different scenarios.
-In the (0,0) case, where all equipment was available, the increase was 18.4%.
-With six trucks offline (0,6), the improvement rose to 22.7%.
-When one shovel was offline (1,0), production volume increased by 14.3%.
-Under both constraints (1,6), RL achieved the highest improvement at 23.3%.
-The RL scheduling (RL-PPO) consistently outperforms random scheduling baseline method across multiple metrics and scenarios, maintaining higher productivity, especially mid-shift.
-Its intelligent decisions optimize resource use, making it highly effective in real-world mining operations with equipment constraints, leading to increased production.
-VII Conclusion and Future Work
-The introduction of Mining-Gym addresses a critical gap in mining process optimization by providing a configurable, open-source benchmarking environment for RL algorithm evaluation in truck dispatch scheduling.
-Our framework offers high-fidelity DES modeling of real-world complexities, seamless integration with OpenAI Gym, and comprehensive visualization tools.
-Preliminary experiments show that RL-based scheduling achieves up to 27.2% higher productivity and 23.3% greater production volume, with the most significant gains in resource-constrained scenarios.
-Mining-Gym provides a standardized evaluation framework, ensuring reproducible research and supporting the adoption of RL solutions in industrial mining operations.
-The development of Mining-Gym as a benchmarking environment will continue in the following directions:
-  • Enhanced Simulation Fidelity: Future iterations will improve traffic modeling, incorporate geographical factors, model heterogeneous equipment fleets, and include additional operational constraints to enhance simulation realism.
-  • Expanded Scenario Library: We will develop a broader set of standardized test scenarios to cover diverse challenges such as weather disruptions, varying ore characteristics, and dynamic production targets.
-  • Multi-objective Optimization Benchmarks: The framework will be extended to support multi-objective optimization, balancing production, cost, equipment lifespan, and environmental impacts—reflecting real-world mining decision-making.
-  • Integration with Digital Twin Capabilities: Enhancing Mining-Gym with digital twin functionalities will connect simulations with real-world data, facilitating data-driven optimizations and bridging the gap between simulation and operational implementation.
-Acknowledgments
-This research was partly supported by the Advance Queensland Industry Research Fellowship AQIRF024-2021RD4.
+Table 8: Distributions for stochastic parameters.
+Parameters are grouped as Sample Once (fixed per environment) and Sample per Episode/Shift (varies daily).
+MTBF: Mean Time Between failures, MTTR: Mean Time To Repair.
+Parameter Group Parameter Distribution / Value
+A. Sample Once (Environment-fixed)
+Equipment counts Trucks (TR) 60
+Shovels (SH) 8
+Crushers (CR) 5
+Dumps (DS) 3
+Equipment characteristics Truck payload (LO) 100 t
+Empty Truck Speed (TE) Normal(55, 12) km/hr
+Loaded Truck Speed (TL) Normal(30, 8) km/hr
+Shovel bucket capacity (SCP) Not specified
+Equipment Maintenance MTBF Shovel (FSH) Poisson(75)
+MTBF Truck (FTR) Poisson(90)
+MTBF Crusher (FCR) Poisson(120)
+MTBF Dumping Site (FDS) Poisson(180)
+MTTR Shovel (RSH) Poisson(60)
+MTTR Truck (RTR) Poisson(45)
+MTTR Crusher (RCR) Poisson(90)
+MTTR Dumping Site (RDS) Poisson(30)
+Infrastructure travel times Shovel → Dump (SZxD) Zone 1: Normal(18, 3)
+Zone 2: Normal(12, 2), Zone 3: Normal(8, 1)
+Shovel → Crusher (SZxC) Zone 1: Normal(8, 2)
+Zone 2: Normal(15, 3), Zone 3: Normal(22, 4)
+B. Sample per Episode / Shift (Varies Daily)
+Fuel consumption Truck with ore (FO) Normal(0.28, 0.05) lt/ton-km
+Truck empty (FE) Normal(0.20, 0.06) lt/ton-km
+Shovel (epsilon) Normal(0.35, 0.07) lt/hr
+Truck Loading Times Zone 1 (TRL_C1) Normal(6, 1) min
+Zone 2 (TRL_C2) Normal(9, 2) min
+Zone 3 (TRL_C3) Normal(12, 2) min
+Truck Dumping / Unloading At Dump (TRDM) Normal(5, 2) min
+At Crusher (TRCR) Normal(15, 1) min
+Experimental Variations Shovels to fail (STF) Uniform(1, 5)
+Trucks to fail (TTF) Uniform(10, 30)
+Shovel initial breakdown (SIB) Uniform(10, 45)
+Truck initial breakdown (TIB) Uniform(10, 90)
+Table 9: Reward function hyperparameters used in demonstration experiments, balancing production efficiency ($\omega_{1}$, $\omega_{2}$), performance bonuses and thresholds ($b_{1}$, $b_{2}$, $\theta_{1}$, $\theta_{2}$, $\phi_{1}$, $\phi_{2}$), and immediate reward parameters ($\alpha$, $\beta$, $\gamma$, $\delta$, $k$, $\lambda$, $|W_{\text{div}}|$, $|W_{\text{streak}}|$).
+All parameters are configurable in Mining-Gym for alternative objectives and ablation studies.
+Function Component Parameter Description Value
+Episodic Reward (Eq. 5) $\omega_{1}$: Production efficiency weight 0.4
+$\omega_{2}$: Diversity penalty weight 0.6
+Immediate Reward Weights (Eq. 3) $\alpha$: Trip time penalty weight 0.25
+$\beta$: Shovel queue time penalty weight 0.15
+$\gamma$: Diversity score penalty weight 0.30
+$\delta$: Streak penalty weight 0.30
+Performance Bonus (Eq. 6) $b_{1}$: High-performance bonus value 0.5
+$b_{2}$: Moderate-performance bonus value 0.2
+Production Thresholds (Eq. 6) $\theta_{1}$: High production threshold 0.95
+$\theta_{2}$: Moderate production threshold 0.90
+Diversity Thresholds (Eq. 6) $\phi_{1}$: High diversity threshold 0.65
+$\phi_{2}$: Moderate diversity threshold 0.50
+Exponentially Weighted Sliding Window (Eq. 4) $k$: Window size (decision points) 5
+$\mu$: Exponential decay rate 0.5
+$n$: Effective window size $\min(k,|D|)$
+Diversity Score Calculation $|W_{\text{div}}|$: Recent decisions window 30
+$\rho$: Numerical stability constant $10^{-10}$
+Streak Penalty Calculation $|W_{\text{streak}}|$: Consecutive decision window 15
+Table 10: Truck Specifications
+Maker Model Payload (MT) Speed (loaded km/hr) Fuel Cons. ($t\textperiodcentered km^{-1}$)
+Komatsu 830E-2 231 59-64 0.3 - 0.4 liters
+CAT 793F 231 55-60 0.35 - 0.45 liters
+Terex MT4400 AC 218 60-64 0.35 - 0.45 liters
+CAT 777(05) 98.2 62 - 67.1 0.25 – 0.35 liters
+Table 11: Shovel and Crusher Specifications
+Model Fuel per Hour Running Cost (USD)
+Shovel Specifications
+CAT 6040/6050 Hydraulic Shovels 250-300 liters per hour \$3.50 to \$4.50 /t
+Komatsu PC8000 946 liters per hour \$4.00 - \$5.00 / t
+Hitachi EX8000-6 1100 liters per hour \$4.50 - \$5.50 / t
+Crusher Specifications
+Metso Lokotrack LT106 1.5 to 2 gallons of diesel per hour \$4.50 to \$6 per hour
+Appendix C Event-driven or decision-point-based RL setting
+The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment.
+This event-driven approach, illustrated in Fig.6, ensures efficient agent-environment interaction aligned with real mining operations, where continuous interaction is impractical.
+The agent observes the current state (including shovel and truck attributes) following the Load-Haul-Dump-Return-Query (LHDRQ) cycle, and makes resource allocation decisions using learned dispatching policies.
+As shown in the figure, the RL policy processes environment states to optimize shovel assignments based on multiple factors, including queue length and equipment status.
+The simulation handles stochastic elements like equipment failures through the Preemption Handler, while the Resource Handler manages resource availability and assignments.
+Figure 6: The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment.